### v2

# STORM: A Spatio-Temporal Factor Model Based on Dual Vector Quantized Variational Autoencoders for Financial Trading

Abstract.
In financial trading, factor models are widely used to price assets and capture excess returns from mispricing. Recently, we have witnessed the rise of variational autoencoder-based latent factor models, which learn latent factors self-adaptively. While these models focus on modeling overall market conditions, they often fail to effectively capture the temporal patterns of individual stocks. Additionally, representing multiple factors as single values simplifies the model but limits its ability to capture complex relationships and dependencies. As a result, the learned factors are of low quality and lack diversity, reducing their effectiveness and robustness across different trading periods. To address these issues, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders, named STORM, which extracts features of stocks from temporal and spatial perspectives, then fuses and aligns these features at the fine-grained and semantic level, and represents the factors as multi-dimensional embeddings. The discrete codebooks cluster similar factor embeddings, ensuring orthogonality and diversity, which helps distinguish between different factors and enables factor selection in financial trading. To show the performance of the proposed factor model, we apply it to two downstream experiments: portfolio management on two stock datasets and individual trading tasks on six specific stocks. The extensive experiments demonstrate STORM’s flexibility in adapting to downstream tasks and superior performance over baseline models. Code is available in PyTorch [^1] https://github.com/DVampire/Storm.
Factor Model, VQVAE, Portfolio Management, Algorithmic Trading
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the ACM Web Conference 2025; April 28–May 02, 2025; Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Computing methodologies Machine learningccs: Applied computing Electronic commerce
1. Introduction
In financial trading, factor models are fundamental tools for asset pricing, widely used to predict asset returns. These models enhance the accuracy of asset pricing and risk management by identifying a set of key factors that explain excess returns. The well-known Fama-French three-factor model (Fama and French, 1992), for example, builds on the Capital Asset Pricing Model (Sharpe, 1964) by introducing additional factors to capture the risk premiums associated with different types of stocks in the market. As financial theories have evolved, researchers have continued to discover new factors, allowing for a more comprehensive understanding of market dynamics.
Figure 1. Overall architecture of STORM
Recently, we have witnessed the rise of latent factor models (Gu et al., 2021; Duan et al., 2022; Wei et al., 2022, 2023), connecting the factor model with the generative model, the variational autoencoder (VAE). These VAE-based models describe high-dimensional data (prices) to low-dimensional representations (factors), and learn factors self-adaptively. Although latent factor models have demonstrated substantial success in financial trading tasks, they still face several significant issues:
  • CH1: Limited Reflection of Market Complexity. Latent factor models represent factors as single values are inherently constrained by their insufficient capacity to capture the intricate complexity and nonlinearity of financial data, rendering them vulnerable to noise and non-stationarity, which compromises their predictive accuracy and stability.
  • CH2: Factor Inefficiency. In addition to the potential limitations of factors discussed in CH1, factors learned through VAEs exhibit three primary inefficiencies: i) Most latent factor models only focus on cross-sectional factors, neglecting the temporal perspective. ii) The continuous, multi-dimensional latent space may allow noise to overshadow meaningful factors. iii) The lack of independence among factors may lead to multicollinearity, with the homogeneity weakening the model’s adaptability to varying market conditions and limiting the effectiveness of factors in downstream investment tasks.
  • CH3: Lack of Factor Selection. Existing latent factor models primarily focus on generating factors without adequately differentiating between them. Furthermore, they neglect the crucial process of factor selection, which is essential for identifying impactful factors, thereby limiting the model’s overall effectiveness and precision.
In order to address the challenges, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders (VQ-VAE), named STORM, with the architecture shown in Figure 1. We represent the factors as multi-dimensional embeddings to capture the complexity and nonlinearity of financial data, thereby improving the factors’ ability to account for factor returns (CH1). Additionally, we develop a dual VQ-VAE architecture to capture cross-sectional and time-series features, considering both spatial and temporal perspectives [^2] We treat cross-sectional as spatial and time-series as temporal. These terms will be used interchangeably unless specified.. We integrate cross-sectional and time-series features at both fine-grained and semantic levels to construct effective factors. To further enhance the model, we introduce diversity loss and orthogonality loss to ensure factor embedding diversity and independence (CH2). Furthermore, we treat the embeddings in the codebook as cluster centers for factor embeddings, effectively using them as class tokens to distinguish between factors. This approach makes the factor differentiation and selection process clear and transparent (CH3). Specifically, our contributions are three-fold:
  • We design a dual VQ-VAE architecture to construct cross-sectional and time-series factors from both spatial and temporal perspectives.
  • We leverage vector quantization techniques to improve factor embedding representation by reducing noise, enhancing diversity, and ensuring orthogonality. This approach significantly strengthens the factors’ predictive power and a more transparent factor selection process.
  • Experiments on two U.S. stock markets across the stock future return prediction task and two downstream tasks (i.e., portfolio management and algorithmic trading) show that STORM outperforms all baselines across 2 prediction metrics and 6 standard financial metrics.
2. Related Work
2.1. Factor Model
Factor model has attracted extensive research focus from the finance community. Financial experts have carefully selected a wide range of factors from macroeconomic (Sharpe, 1964), market changes (Fama and French, 1992), individual stock fundamentals (Greenblatt, 2010), and other perspectives to price assets in an attempt to capture excess returns from mispricing. Recently, latent factor models have emerged that improve investment returns by studying latent factors self-adaptively. Specifically, (Gu et al., 2021) learns latent factors depend on assets characteristics, (Duan et al., 2022) learns cross-sectional factors by predicting future returns, and (Wei et al., 2023) learns market-augmented latent factors. However, these VAE-based latent factor models may suffer from a lack of robustness in representing low signal-to-noise ratio markets, leading to reduced reconstruction and generation capabilities. Additionally, their effectiveness is constrained by a primary focus on with insufficient attention to the temporal perspective.
2.2. Financial Trading
Financial trading has attracted significant attention from both the finance and AI communities. Quantitative traders leverage mathematical models and algorithms to automatically identify trading opportunities (Sun et al., 2024). Within the Fin-tech field, research tasks include portfolio management (PM), algorithmic trading (AT), order execution (Fang et al., 2021), market making (Spooner and Savani, 2020), etc.
PM focuses on optimizing wealth allocation across assets (Li and Hoi, 2014). Its development has evolved from simple rule-based methods (Poterba and Summers, 1988) to prediction-based approaches using machine learning (Ke et al., 2017) and deep learning (Qin et al., 2017; Zhao et al., 2022) for predicting returns or price movements. Despite recent integration of reinforcement learning (RL) in PM (Zhang et al., 2024b), the latent factor models still rely on deep learning for market modeling and precise predictions.
AT involves trading financial assets using algorithm-generated signals. Similar to PM, there are rule-based and supervised learning methods (Yang et al., 2020) in the AT task. In recent years, RL-based methods (Sun et al., 2022) have become dominant for their superiority in complex sequential decision-making.
2.3. Vector Quantized Variational Autoencoder
VQ-VAE has made significant contributions across several areas of deep learning, particularly in representation learning, generative modeling, and time series analysis. Initially introduced by Van Den Oord et al., VQ-VAE has showcased its exceptional capabilities in processing high-dimensional data, e.g., images (Razavi et al., 2019), audio (Zhang et al., 2024a), and video (Yan et al., 2021), with a strong emphasis on learning discrete latent representations. Moreover, VQ-VAE has proven highly effective in handling complex time-dependent data (Talukder et al., 2024), advancing the development of robust generative models specifically in the context of temporal analysis.
3. Preliminaries
3.1. Factor Model
The factor model (e.g., Arbitrage Pricing Theory (Ross, 1976)) is defined to study the relationship between expected asset returns and factor exposures, focusing on the cross-sectional differences in expected asset returns.
(1) $\mathbb{E}[R_{i}^{e}]=\alpha_{i}+\beta_{i}\lambda\text{.}$
$\mathbb{E}[\cdot]$ denotes the expectation operator. $\mathbb{E}[R_{i}^{e}]$ is the expected excess return. $\beta_{i}$ is factor exposure (or factor loading) of asset $i$, and $\lambda$ is factor expected return (or factor risk premium). $\alpha_{i}$ is pricing error.
According to (Duan et al., 2022), they define the factor model in the spatial perspective, which only focuses on why the expected returns of different assets vary, rather than how the returns of each asset change over time. Therefore, we expand it from the temporal perspective and introduce the general functional form of the latent factor model:
(2) $y_{i,t}=\alpha_{i,t}+\sum_{k=1}^{K}\bm{\beta}_{i,t}^{k}\mathbf{z}_{i,t}^{k}+\epsilon_{i,t}\text{,}\vspace{-0.1cm}$
where $y_{i,t}$ denotes the future return of stock $i$ at timestep $t$, $\mathbf{z}_{i,t}$ denotes $K$ factor returns, and $\epsilon_{i,t}$ is the noise. From the spatial, Equation 2 can be expressed as $\mathbf{y}_{i}=\bm{\alpha}_{i}+\sum_{k=1}^{K}\bm{\beta}_{i}^{k}\mathbf{z}_{i}^{k}+\bm{\epsilon}_{i}$, which we regard as the cross-sectional factor model. From the temporal perspective, Equation 2 can be expressed as $\mathbf{y}_{t}=\bm{\alpha}_{t}+\sum_{k=1}^{K}\bm{\beta}_{t}^{k}\mathbf{z}_{t}^{k}+\bm{\epsilon}_{t}$, which we regard as the time-series factor model. The cross-sectional factor model focuses on the differences in returns of different assets at the same point in time, while the time-series factor model focuses on the changes in returns of a single asset at different points in time.
3.2. Problem Formulation
To assess the effectiveness and generating ability of factors, we predict and evaluate the stock future returns. To show the adaptability of the latent factor model with multiple downstream tasks, we employ two financial trading tasks, i.e., portfolio management (Wang et al., 2021; Zhang et al., 2024b; Ye et al., 2020) and algorithmic trading (Sun et al., 2022; Liu et al., 2020), to showcase the investment capabilities.
Observed Data: we utilize the stock’s historical price data $\mathbf{p}$(i.e., open, high, low, and close) and technical indicators $\mathbf{d}$ as observed variables $\mathbf{x}:=[\mathbf{p},\mathbf{d}]\in\mathbb{R}^{N\times W\times D}$, within a window size of $W$. The historical price data of stock $i$ at time $t$ is denoted as $\mathbf{p}_{i,t}:=[p_{i,t}^{o},p_{i,t}^{h},p_{i,t}^{l},p_{i,t}^{c}]\in\mathbb{R}^{D_{1}}$. The technical indicators which are calculated from price and volume values, are denoted as $\mathbf{d}_{i,t}\in\mathbb{R}^{D_{2}}$. Each stock is represented by $D=D_{1}+D_{2}$ features per trading day, which we then use to predict the stock’s future returns, defined as $\mathbf{y}_{i,t+1}:=\frac{p_{i,t+1}-p_{i,t}}{p_{i,t}}$.
Downstream tasks:
  • Portfolio management task aims to construct a portfolio based on the predicted returns to test the profitability of the factor model. At each trading day $t$, we obtain the portfolio weight $\mathbf{w}_{t}=[w_{t}^{0},w_{t}^{1},...,w_{t}^{N-1}]\in\mathbb{R}^{N}$ for $N$ stocks. Each stock $i$ is assigned a weight $w_{t}^{i}$ representing its portfolio proportion, subject to the constraint that $\sum_{i=0}^{N-1}w_{t}^{i}=1$ for full investment.
  • Algorithmic trading task aims to execute buy, hold, and sell actions based on predicted asset states to balance returns and risks. We formulate it as a Markov Decision Process (MDP) under reinforcement learning scenario following (Chen et al., 2023; Zhang et al., 2024c). The MDP is constructed by a 5-tuple $(\mathit{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where $\mathit{S}$ and $\mathcal{A}$ are sets of states and actions respectively, $\mathcal{T}:\mathit{S}\times\mathcal{A}\times\mathit{S}\rightarrow[0,1]$ is the state transition function, $\mathcal{R}:\mathit{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function where $\mathbb{R}$ is a continuous set of possible rewards, and $\gamma\in[0,1)$ is the discount factor. The goal is to learn a policy $\mathcal{\pi}:\mathit{S}\rightarrow\mathcal{A}$ that maximizes the expected discounted cumulative reward $\mathbb{E}[\sum^{\infty}_{t=0}\gamma^{t-1}r_{t}]$. At each trading day $t$ for specific stock $i$, the agent takes action $a_{i,t}:=\{a_{i,t}^{b},a_{i,t}^{h},a_{i,t}^{s}\}\in\mathcal{A}$ which means buy, hold and sell, according to the current environment $s_{t}\in\mathcal{S}$.
3.3. Vector Quantized Variational Autoencoder
VQ-VAE(Van Den Oord et al., 2017) is different from VAE (Kingma and Welling, 2013) as it quantizes the observed data into a discrete token sequence. The VQ-VAE contains three parts, encoder, decoder, and discrete codebook, denoted by $\phi_{\text{enc}}$ with parameter $\theta_{\text{enc}}$, $\phi_{\text{dec}}$ with parameter $\theta_{\text{dec}}$ and $\mathcal{C}=\{(k,e_{k}\in R^{H})\}_{k=1}^{K}$ with parameter $\{e_{k}\in R^{H}\}_{k=1}^{K}$ respectively. Among them, $K$ is the size of the codebook, $H$ is the dimension of the vector, and $e_{k}$ is the $k$-th vector of the latent embedding space. Given an input $x$, the encoder $\phi_{\text{enc}}$ firstly encodes it into a set of continuous feature vectors $Z=\phi_{\text{enc}}(x)$ and quantization operation $q(\cdot)$ passes each feature vector $z\in Z$ through the discretization bottleneck following by selecting the nearest neighbor embedding in the codebook $\mathcal{C}$ as its discrete code sequence $Q_{z}$.
(3) $Q_{z}=q(z,\mathcal{C})=c,\text{where }c=\arg\min_{k\in[0,K-1]}||z-e_{k}||_{2}^{2}\text{.}\vspace{-0.1cm}$
Then, the quantized feature for $z$ denoted as $Z^{q}_{z}$, is obtained through $Z^{q}_{z}=e_{Q_{z}}$. The quantized vectors $Z^{q}$ is fed into the decoder $\theta_{\text{dec}}$ to reconstruct the original data $x^{\prime}=\theta_{\text{dec}}(Z^{q})$.
Figure 2. The architecture of our proposed STORM model
4. Method
In this section, we represent our proposed latent factor model STORM, with the overall structure of the VQ-VAE depicted in Figure 2. In general, STORM models data into time-series and cross-sectional features, deriving factors by integrating both temporal and spatial perspectives. Firstly, we introduce the dual VQ-VAE architecture and explain the extraction of features in two modules. Then the factor module will fuse and align features and generate factor embeddings. Finally, we demonstrate the effectiveness of STORM through comprehensive evaluations on two critical downstream tasks: portfolio management and algorithmic trading.
4.1. Dual VQ-VAE Structure
The dual VQ-VAE architecture processes stock data from spatial and temporal perspectives by patching the data in both modules. Then, the encoder extracts time-series (TS) and cross-sectional (CS) features, which are subsequently clustered using a learnable codebook for quantized representation. Finally, the decoder reconstructs the data, capturing complex patterns and enhancing forecasting performance.
4.1.1. Patching
In stock data analysis, it is essential to extract features from both temporal and spatial dimensions. Time-series features capture the dynamics of a stock over time, while cross-sectional features reveal the correlations among different stocks. To achieve this, we partition data into patches (Nie et al., 2022) within the TS and CS modules. In the TS module, the observed data $\mathbf{x}$ is divided along the stock number dimension, with each patch containing data for one stock over $p$ days. Similarly, in the CS module, the observed data $\mathbf{x}$ is divided along the time axis, with each patch containing features of all stocks for a single trading day. Combining information from multi-scale features allows the model to capture both temporal and spatial dependencies, enhancing the depth and accuracy of market insights.
4.1.2. TS and CS Encoders
Considering the Transformer’s (Vaswani et al., 2017) remarkable ability to capture both local and global dependencies via self-attention and patching strategies, it is particularly well-suited for long-term time-series forecasting (Nie et al., 2022). Therefore, we employ stacked transformer blocks as the encoders in the TS and CS modules to capture complex time-series and cross-sectional patterns, ultimately improving the accuracy and robustness of feature extraction.
4.1.3. Codebook Construction and Optimization
The construction of the codebook involves a learnable embedding space designed to quantize the encoded features into discrete representations (Van Den Oord et al., 2017). During training, the continuous latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$ are quantized into discrete space by mapping each vector to its nearest codebook entry, $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$, based on minimizing the Euclidean distance. This process translates the encoded features into discrete tokens, enabling the model to leverage a finite set of vectors to capture complex patterns.
The VQ-VAE offers significant advantages in the latent factor model, primarily through its use of discrete codebook embeddings: i) Discretization Benefits: The discretization process inherent in VQ-VAE helps in clustering distinct and meaningful factors, providing orthogonality and diversity which improves upon the factors derived from traditional methods like FactorVAE. ii) Explicit Factor Selection: The discrete token indices facilitate an explicit factor selection process, identifying the most relevant factors that influence stock returns and thus enhancing prediction accuracy. iii) Noise Reduction: The factor selection process filters out irrelevant or redundant information, reducing noise and improving the overall robustness of the model.
To encourage equal usage of the codebook vectors (Baevski et al., 2020), we incorporate a diversity loss to ensure balanced representations across the codebook. It is achieved by maximizing the entropy of the averaged softmax distribution over vectors for each codebook $\bar{p}_{g}$:
(4) $\mathcal{L}_{\text{div}}=\frac{1}{GK}\sum_{g=1}^{G}\sum_{k=1}^{K}\bar{p}_{g,k}\log{\bar{p}_{g,k}}\text{,}\vspace{-0.2cm}$
where $G$ represents the number of codebooks, set 2 in STORM. The diversity constraint helps enhance representational capacity, leading to better coverage of the input space.
However, the lack of independence among factors can lead to multicollinearity, causing the model to overfit and are less robust to various market situations. To address this issue, we impose an orthogonality constraint on the codebook, as recent studies (Shin et al., 2023) have shown that enforcing orthogonality allows discretized codes to maintain translation equivariance. The orthogonality loss is:
(5) $\mathcal{L}_{\text{ortho}}=\frac{1}{K^{2}}\left|\left|\ell_{2}(\mathbf{e})^{\top}\ell_{2}(\mathbf{e})-I_{K}\right|\right|_{F}^{2}\text{,}\vspace{-0.1cm}$
where $I_{k}\in\mathbb{R}^{K\times K}$ is the identity matrix, $\mathbf{e}$ is a surrogate for two codebook embeddings $\mathbf{ts}$ and $\mathbf{cs}$, $\ell_{2}(\mathbf{e})$ denotes $L_{2}$-normalized embeddings, and $\left|\left|\cdot\right|\right|_{F}$ is the Frobenius norm. The orthogonal constraint ensures that factors are independent of each other, allowing the effect of each factor on returns to be individually explained, which is crucial for asset pricing analysis.
4.2. TS and CS Decoders
The decoder is responsible for reconstructing the original data $\mathbf{x}$ from the quantized latent vectors $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$. We utilize Transformer as decoders either, to generate constructed data $\mathbf{x}_{ts}^{\prime}$ and $\mathbf{x}_{cs}^{\prime}$, which aim to closely approximate the original input data.
Finally, the encoders, decoders, and codebooks are jointly learned by minimizing the following loss objectives:
(6) $\mathcal{L}_{1}=$ $\lambda_{\text{ortho}}\mathcal{L}_{\text{ortho}}+\lambda_{\text{div}}\mathcal{L}_{\text{div}}+\left|\left|\mathbf{x}-\mathbf{x}^{\prime}_{ts}\right|\right|^{2}_{2}+\left|\left|\mathbf{x}-\mathbf{x}^{\prime}_{cs}\right|\right|^{2}_{2}$
$+$ $\left|\left|sg[\mathbf{z}_{e}^{ts}(\mathbf{x})]-\mathbf{z}_{q}^{ts}(\mathbf{x})\right|\right|^{2}_{2}+\left|\left|sg[\mathbf{z}_{q}^{ts}(\mathbf{x})]-\mathbf{z}_{e}^{ts}(\mathbf{x})\right|\right|^{2}_{2}$
$+$ $\left|\left|sg[\mathbf{z}_{e}^{cs}(\mathbf{x})]-\mathbf{z}_{q}^{cs}(\mathbf{x})\right|\right|^{2}_{2}+\left|\left|sg[\mathbf{z}_{q}^{cs}(\mathbf{x})]-\mathbf{z}_{e}^{cs}(\mathbf{x})\right|\right|^{2}_{2}\text{,}$
where $sg[\cdot]$ is a stop-gradient operator.
4.3. Factor Module
The factor module aims to integrate the TS and CS latent features through feature fusion and alignment. Then, through prior-posterior learning to generate factor embeddings from latent features to predict stock future returns.
4.3.1. Feature Fusion and Alignment
As mentioned above we employ different patching methods to ensure that both time-series and cross-sectional features of the data were accurately captured. However, these methods result in two parallel processes that do not align at the feature level. To address this limitation, we employ a cross-attention mechanism to enhance the interaction and fusion of features at a fine-grained level, improving the model’s dynamic understanding of the input data. Additionally, contrastive learning is used at the semantic level to enhance feature representation by emphasizing semantic similarities and differences.
Cross-Attention Mechanism. In order to fuse the features in the TS and CS modules at a fine-grained level, we leverage a multi-scale encoder to combine information between two patches and produce stronger fusion features for factor generations, motivated by CrossViT (Chen et al., 2021). The encoder contains cross-attention layers that achieve both accuracy and efficiency.
Contrastive Learning. We incorporate contrastive learning, inspired by its success in aligning text and image features through a shared embedding space (Li et al., 2023; Radford et al., 2021), to enhance semantic consistency between time-series and cross-sectional features. In the factor module, contrastive loss ensures that similar features cluster together while dissimilar features are pushed apart in the embedding space.
4.3.2. Prior-Posterior Learning
Through multi-scale encoder and contrastive learning layer, the CS and TS latent features are fused and aligned. In order to retain the codebook’s categorical properties for factors, we add the codebook embeddings $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$ as extra [CLS]-tokens (Dosovitskiy et al., 2020) to the latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$, respectively. Then, two latent features concat together to get the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$.
Due to stock market’s inherent volatility and complexit, it is difficult to bridge the gap between the noisy market and effective factor model (Duan et al., 2022). Therefore, we utilize the prior-posterior structure to predict future returns and optimize latent factor embeddings. In the training stage, the posterior layer predict posterior distribution of factor expected returns $\mathbf{z}_{post}$ from the true future stock returns $\mathbf{y}$ and the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$:
(7) $[\mu_{post},\sigma_{post}]=\phi_{FE}(\mathbf{y},\mathbf{z}_{e}(\mathbf{x}))\text{,}$
and $\mathbf{z}_{post}$ follows the independent Gaussian distribution. In the inference stage, the prior layer only use latent factor embeddings without any future information leakage to calculate prior distribution of factor expected returns $\mathbf{z}_{prior}$:
(8) $[\mu_{prior},\sigma_{prior}]=\phi_{FP}(\mathbf{z}_{e}(\mathbf{x}))\text{.}$
The return predictor uses factor expected return $\mathbf{z}$ to calculate future returns:
(9) $\hat{\mathbf{y}}=\bm{\alpha}+\sum_{k=1}^{K}\bm{\beta}^{k}\mathbf{z}^{k}+\bm{\epsilon}\text{.}$
4.4. Downstream Tasks
Existing research on latent factor models (Duan et al., 2022; Wei et al., 2023) primarily focuses on predicting future returns by analyzing cross-sectional characteristics of market stocks, which means the profitability of the factor model in quantitative trading is typically evaluated through portfolio management tasks. In contrast, STORM captures both the cross-sectional features among stocks and the time-series features of individual stocks, enabling it to excel in both PM and single-asset AT tasks.
4.4.1. Portfolio Management
To evaluate the performance of the STORM and compare it with baseline methods, we follow the paradigm proposed by FactorVAE (Duan et al., 2022), focusing on return prediction and the PM task. Specifically, we utilize the factor decoder network to generate stock future returns $\mathbf{\hat{y}}$, and then apply the TopK-Drop strategy [^3] https://qlib.readthedocs.io/en/latest/component/strategy.html to backtest the factor model. The TopK-Drop strategy constructs a daily portfolio consisting of top $k$ stocks based on the return predictions. Considering the transaction costs in the real stock market, the strategy sets the turnover constraint, which ensures that more than $k-d$ stocks overlap between the portfolios $\mathbf{w}_{t}$ and $\mathbf{w}_{t+1}$ on consecutive trading days.
4.4.2. Algorithmic trading
In STORM, $\mathbf{z}_{e}(\mathbf{x})$ is the stock latent factor embeddings encoded by the dual VQ-VAE architecture, including both time-series and cross-sectional factor embeddings in the data. The latent factor embeddings $\mathbf{Z}$ are integrated into the observation set $\mathcal{O}=\{\mathbf{Z},\mathcal{R}\}$, where $\mathcal{R}$ is the reward function used to guide the agent’s learning and decision-making in the environment. By combining latent factor embeddings with reward, the agent can more accurately understand the market environment and adjust its strategy based on observations. In the AT task, since it involves trading a single stock, the features of other stocks are excluded. We use the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) to optimize the policy.
5. Experiment
In this section, we evaluate the proposed STORM on real stock markets and conduct extensive experiments to address the following research questions. RQ1: How to evaluate the effectiveness of STORM’s learned factors? RQ2: How does STORM perform on downstream tasks? RQ3: How do the key components contribute to the performance of STORM?
Models SP500 DJ30
[…]
Future Return Prediction in Portfolio Management Task
LightGBM 0.027 0.274 0.031 0.272
± 0.006 ± 0.084 ± 0.005 ± 0.049
LSTM 0.034 0.333 0.031 0.329
± 0.006 ± 0.042 ± 0.004 ± 0.056
Transformer 0.035 0.340 0.033 0.343
± 0.007 ± 0.078 ± 0.005 ± 0.045
CAFactor 0.037 0.356 0.040 0.380
± 0.005 ± 0.084 ± 0.003 ± 0.043
FactorVAE 0.052 0.543 0.056 0.520
± 0.010 ± 0.122 ± 0.012 ± 0.081
HireVAE 0.057 0.558 0.058 0.563
± 0.006 ± 0.058 ± 0.006 ± 0.053
STORM 0.063 0.679 0.066 0.670
± 0.020 ± 0.169 ± 0.038 ± 0.303
STORM-w/o-TS 0.051 0.514 0.055 0.560
± 0.016 ± 0.141 ± 0.016 ± 0.105
STORM-w/o-CS 0.053 0.514 0.053 0.586
± 0.013 ± 0.126 ± 0.017 ± 0.143
Improvement(%)11footnotemark: 1 10.526 21.685 13.793 19.005
  • $1$
    Improvement of STORM over the best-performing baselines.
Table 1. Results of all models on two metrics, i.e., RankIC and RankICIR (mean ± range, computed across 10 runs).
5.1. Experiment Settings
Dataset. We conduct experiments on two U.S. stock markets, SP500 and DJ30, using stock daily data that includes technical features based on Alpha158 (Yang et al., 2020). Both datasets span 16 years, from 2008-04-01 to 2024-03-31, encompassing global conditions, e.g., the 2007-2008 financial crisis and COVID-19. Datasets are chronologically divided into non-overlapping training (from 2008-04-01 to 2021-03-31) and test (from 2021-04-01 to 2024-03-31) sets.
Metrics. We compare STORM and baselines in terms of 6 financial metrics across the PM and AT tasks. The financial metrics include 2 profit criteria, Annualized Percentage Yield (APY) and Cumulative Wealth (CW), 2 risk-adjusted profit criteria, Calmar Ratio (CR) and Annualized Sharpe Ratio (ASR), and 2 risk criteria, Maximum Drawdown (MDD) and Annualized Volatility (AVO). The calculation formulas and meanings of these metrics are as follows:
  • CW is the total returns yielded from a portfolio strategy: $\text{CW}_{T}=\prod_{i=1}^{T}(1+r_{i})$, where $r_{i}$ is the net return.
  • APY measures the average wealth increment that one portfolio strategy could achieve compounded in a year, which is defined as $\text{APY}_{T}=\sqrt[y]{\text{CW}_{T}}-1$, where $y$ is the number of years corresponding to $T$ trading rounds.
  • MDD measures the largest loss from a historical peak in the cumulative wealth to show the worst case, which is defined as: $MDD=\mathop{\max}_{i=0}^{T}\frac{P_{i}-R_{i}}{P_{i}}$, where $R_{i}=\prod_{i=1}^{T}{\frac{V_{i}}{V_{i-1}}}$ and $P_{i}=\mathop{\max}_{i=1}^{T}R_{i}$.
  • AVO is the annualized standard deviation of daily returns and multiplied by $\sqrt{AT}$, where $AT$ is the average trading rounds of annual trading days and $AT=252$ for all the datasets.
  • ASR is an annualized volatility risk-adjusted return defined as $\text{ASR}=\frac{\text{APY}-R_{f}}{\text{AVO}}$, where $R_{f}$ is the risk-free return.
  • CR measures the drawdown risk-adjusted return of a portfolio calculated as $CR=\frac{\text{APY}}{\text{MDD}}$.
Typically, to evaluate the effectiveness of the learned factors in the PM task, we adopt the Rank Information Coefficient (RankIC) and the Information Ratio of RankIC (RankICIR):
  • RankIC is a ranking metric in finance, which measures the correlation between the predicted rankings and the actual returns. It is defined as:
    (10) $\text{RankIC}_{s}=\frac{1}{N}\frac{(r_{\hat{y}_{s}}-\text{mean}(r_{\hat{y}_{s}}))^{T}(r_{y_{t}}-\text{mean}(r_{y_{s}}))}{\text{std}(r_{\hat{y}_{s}})\cdot\text{std}(r_{y_{s}})}$
    $\text{RankIC}=\frac{1}{T_{\text{test}}}\sum_{s=1}^{T_{\text{test}}}\text{RankIC}_{s}\text{,}$
    where $T_{test}$ is the number of trading days of the test range, $r_{y_{s}}$ and $r_{\hat{y}_{s}}$ are the true and predicted ranks of stocks on the trading day $s$.
  • RankICIR is the information ratio of RankIC, which measures the stability of prediction, $\text{RankICIR}=\frac{\text{mean}(\text{RankIC}_{s})}{\text{std}(\text{RankIC}_{s})}$.
Baselines. We compare STORM with nine methods across one prediction task and two downstream tasks. The baseline methods can be divided into three categories: ML & DL-based models, Factor models, and RL-based models. In the prediction task and PM task, we compare STORM with ML & DL-based and factor models. In the AT task, we compare STORM with ML & DL-based and RL-based methods. We use the labels P, PM, and AT to indicate the applicable task scope for the following baselines. The following will provide a brief introduction to each method:
  • Market
    • Buy-and-Hold (B&H) involves holding assets for an extended period, regardless of short-term market fluctuations, assuming that long-term returns will be more favorable. (P, PM, AT)
  • ML&DL-based
    • LGBM (Yang et al., 2020) uses a series of tree models to predict price fluctuations and provide buy and sell signals. (P, PM)
    • LSTM (Yu and Li, 2018) utilizes long short-term memory to improve the accuracy of price predictions. (P, PM)
    • Transformer (Yang et al., 2020) models leverage self-attention mechanisms to enhance the precision of price forecasts. (P, PM)
  • RL-based
    • SAC (Haarnoja et al., 2018) is an off-policy actor-critic algorithm that optimizes trading strategies using entropy regularization and soft value functions in continuous action spaces. (PM, AT)
    • PPO (Schulman et al., 2017) updates trading policies iteratively to balance exploration and exploitation, ensuring stability and sample efficiency. (PM, AT)
    • DQN (Mnih et al., 2013) uses deep neural networks to approximate the action-value function and make trading decisions from market data. (PM, AT)
  • Factor-model
    • CAFactor is a state-of-the-art model that learns latent factors depending on asset characteristics. (P, PM)
    • FactorVAE (Duan et al., 2022) is a state-of-the-art VAE-based latent factor model that learns optimal latent stock factors. (P, PM)
    • HireVAE (Wei et al., 2023) is a state-of-the-art VAE-based latent factor model that identifies factors between market situation and stock-wise. (P, PM)
SP500 Dataset
Strategies Profit Risk-Adj. Profit Risk
[…]
Market Index 0.058 1.184 0.228 0.142 0.254 0.410
LightGBM 0.059 1.201 0.304 0.332 0.238 0.176
± 0.132 ± 0.456 ± 0.753 ± 0.755 ± 0.108 ± 0.008
LSTM 0.069 1.221 0.278 0.371 0.248 0.186
± 0.020 ± 0.068 ± 0.065 ± 0.114 ± 0.042 ± 0.013
Transformer 0.076 1.244 0.389 0.433 0.198 0.174
± 0.028 ± 0.098 ± 0.165 ± 0.161 ± 0.057 ± 0.004
CAFactor 0.075 1.241 0.342 0.428 0.223 0.174
± 0.028 ± 0.100 ± 0.229 ± 0.165 ± 0.043 ± 0.006
FactorVAE 0.079 1.256 0.404 0.460 0.200 0.173
± 0.025 ± 0.085 ± 0.177 ± 0.145 ± 0.040 ± 0.007
HireVAE 0.077 1.249 0.361 0.448 0.216 0.172
± 0.029 ± 0.104 ± 0.180 ± 0.189 ± 0.048 ± 0.007
STORM 0.177 1.641 1.193 0.817 0.692 0.072
± 0.103 ± 0.457 ± 0.280 ± 0.071 ± 0.091 ± 0.011
STORM-w/o-TS 0.084 1.278 1.055 0.702 0.632 0.068
± 0.060 ± 0.204 ± 0.271 ± 0.062 ± 0.107 ± 0.003
STORM-w/o-CS 0.080 1.264 0.971 0.698 0.684 0.068
± 0.052 ± 0.176 ± 0.192 ± 0.065 ± 0.071 ± 0.005
Improvement(%) 124.05 30.65 195.30 77.61 - 60.47
DJ30 Dataset
Strategies Profit Risk-Adj. Profit Risk
[…]
Market Index 0.063 1.201 0.147 0.429 0.219 0.288
LightGBM 0.069 1.221 0.288 0.430 0.244 0.160
± 0.040 ± 0.140 ± 0.262 ± 0.267 ± 0.047 ± 0.007
LSTM 0.060 1.192 0.243 0.370 0.248 0.163
± 0.007 ± 0.023 ± 0.021 ± 0.043 ± 0.006 ± 0.005
Transformer 0.056 1.179 0.227 0.367 0.250 0.154
± 0.013 ± 0.043 ± 0.067 ± 0.085 ± 0.033 ± 0.002
CAFactor 0.059 1.186 0.233 0.38 0.252 0.153
± 0.015 ± 0.050 ± 0.058 ± 0.092 ± 0.021 ± 0.001
FactorVAE 0.076 1.246 0.352 0.480 0.225 0.159
± 0.038 ± 0.135 ± 0.349 ± 0.235 ± 0.063 ± 0.005
HireVAE 0.072 1.233 0.298 0.445 0.247 0.163
± 0.037 ± 0.132 ± 0.253 ± 0.256 ± 0.057 ± 0.012
STORM 0.143 1.501 0.839 0.680 0.644 0.057
± 0.078 ± 0.321 ± 0.213 ± 0.080 ± 0.102 ± 0.004
STORM-w/o-TS 0.071 1.244 0.757 0.596 0.634 0.056
± 0.134 ± 0.509 ± 0.375 ± 0.144 ± 0.159 ± 0.004
STORM-w/o-CS 0.077 1.266 0.764 0.607 0.632 0.057
± 0.156 ± 0.612 ± 0.203 ± 0.18 ± 0.118 ± 0.004
Improvement(%) 88.16 20.47 138.35 41.67 - 63.40
Table 2. Portfolio management task results of all models across six metrics (mean ± range, computed across 10 runs).
Models AAPL JPM IBM INTC MSFT
[…]
Buy&Hold 0.120 1.404 0.383 0.096 1.316 0.236 0.145 1.499 0.727 -0.117 0.690 -0.184 0.214 1.784 0.569
LightGBM 0.135 1.390 0.487 0.116 1.335 0.333 0.227 1.654 1.091 -0.042 0.880 0.038 0.267 2.068 0.637
LSTM 0.053 1.152 0.283 0.079 1.290 0.266 0.134 1.386 0.754 0.060 1.262 0.381 0.178 1.513 0.893
± 0.015 ± 0.042 ± 0.101 ± 0.019 ± 0.053 ± 0.023 ± 0.141 ± 0.406 ± 0.69 ± 0.119 ± 0.429 ± 0.62 ± 0.122 ± 0.352 ± 0.822
Transformer 0.083 1.240 0.512 0.133 1.384 0.614 0.131 1.377 0.782 0.079 1.290 0.458 0.138 1.397 0.726
± 0.099 ± 0.284 ± 0.864 ± 0.07 ± 0.201 ± 0.509 ± 0.072 ± 0.207 ± 0.541 ± 0.174 ± 0.68 ± 0.824 ± 0.118 ± 0.338 ± 0.643
DQN 0.135 1.374 0.510 0.105 1.305 0.607 0.139 1.400 0.802 0.061 1.185 0.442 0.166 1.475 0.534
± 0.075 ± 0.226 ± 0.239 ± 0.132 ± 0.383 ± 0.673 ± 0.055 ± 0.158 ± 0.270 ± 0.106 ± 0.305 ± 0.436 ± 0.035 ± 0.100 ± 0.107
SAC 0.147 1.509 0.528 0.131 1.383 0.400 0.207 1.598 1.170 0.056 1.165 0.353 0.229 1.656 0.929
± 0.021 ± 0.060 ± 0.052 ± 0.038 ± 0.111 ± 0.198 ± 0.057 ± 0.163 ± 0.335 ± 0.210 ± 0.6 ± 0.994 ± 0.114 ± 0.331 ± 0.585
PPO 0.137 1.379 0.496 0.128 1.372 0.356 0.146 1.422 0.779 -0.019 0.954 0.040 0.216 1.620 0.569
± 0.073 ± 0.209 ± 0.213 ± 0.026 ± 0.071 ± 0.059 ± 0.047 ± 0.136 ± 0.205 ± 0.177 ± 0.514 ± 0.418 ± 0.081 ± 0.232 ± 0.105
STORM 0.137 1.472 0.501 0.134 1.459 0.363 0.239 1.903 1.263 0.099 1.339 0.433 0.250 1.954 0.724
± 0.015 ± 0.058 ± 0.035 ± 0.039 ± 0.151 ± 0.072 ± 0.033 ± 0.150 ± 0.224 ± 0.085 ± 0.321 ± 0.289 ± 0.034 ± 0.163 ± 0.086
STORM-w/o-TS 0.124 1.428 0.714 0.127 1.437 0.627 0.126 1.428 0.703 -0.001 1.000 0.099 0.103 1.341 0.360
± 0.062 ± 0.229 ± 0.003 ± 0.058 ± 0.214 ± 0.187 0± 0.022 ± 0.085 ± 0.091 ± 0.049 ± 0.152 ± 0.171 ± 0.017 ± 0.062 ± 0.065
STORM-w/o-CS 0.059 1.187 0.309 0.103 1.347 0.350 0.149 1.525 0.908 -0.031 0.913 0.048 0.148 1.513 0.445
± 0.021 ± 0.070 ± 0.061 ± 0.043 ± 0.154 ± 0.106 ± 0.095 ± 0.400 ± 0.450 ± 0.053 ± 0.156 ± 0.109 ± 0.03 ± 0.117 ± 0.040
Improvement(%) -6.803 -2.452 35.227 0.752 5.419 2.117 5.286 15.054 7.949 25.316 3.798 -5.459 -6.367 -5.467 -22.067
Models AAPL JPM IBM INTC MSFT
[…]
Buy&Hold 0.447 0.313 0.269 0.405 0.406 0.237 0.679 0.199 0.213 -0.324 0.635 0.359 0.777 0.376 0.275
LightGBM 0.950 0.309 0.017 0.921 0.386 0.015 1.822 0.175 0.011 0.098 0.553 0.023 1.456 0.370 0.017
LSTM 0.623 0.244 0.012 0.801 0.335 0.012 1.268 0.163 0.010 0.287 0.149 0.010 1.345 0.246 0.014
± 0.163 ± 0.072 ± 0.003 ± 0.185 ± 0.061 ± 0.002 ± 0.924 ± 0.017 ± 0.001 ± 1.02 ± 0.101 ± 0.007 ± 0.842 ± 0.124 ± 0.003
Transformer 0.972 0.239 0.010 1.332 0.226 0.010 1.336 0.182 0.010 0.604 0.257 0.014 1.206 0.182 0.011
± 0.748 ± 0.187 ± 0.006 ± 0.607 ± 0.100 ± 0.003 ± 0.688 ± 0.100 ± 0.003 ± 0.789 ± 0.194 ± 0.006 ± 0.754 ± 0.134 ± 0.006
DQN 0.974 0.288 0.016 1.085 0.212 0.010 1.415 0.161 0.009 0.609 0.243 0.014 1.194 0.309 0.015
± 0.472 ± 0.159 ± 0.007 ± 1.270 ± 0.118 ± 0.003 ± 0.403 ± 0.049 ± 0.004 ± 0.584 ± 0.310 ± 0.009 ± 0.148 ± 0.079 ± 0.004
SAC 1.054 0.292 0.016 1.044 0.350 0.014 1.794 0.157 0.011 0.527 0.359 0.017 1.672 0.229 0.013
± 0.223 ± 0.022 ± 0.002 ± 0.276 ± 0.085 ± 0.001 ± 0.458 ± 0.068 ± 0.005 ± 1.061 ± 0.191 ± 0.004 ± 0.736 ± 0.215 ± 0.005
PPO 0.974 0.300 0.016 1.000 0.381 0.015 1.386 0.177 0.011 -0.044 0.511 0.019 1.276 0.353 0.017
± 0.412 ± 0.030 ± 0.001 ± 0.189 ± 0.018 ± 0.001 ± 0.231 ± 0.027 ± 0.003 ± 1.689 ± 0.115 ± 0.007 ± 0.299 ± 0.036 ± 0.001
STORM 0.913 0.309 0.017 0.983 0.389 0.015 1.843 0.162 0.011 0.700 0.326 0.019 1.441 0.312 0.016
± 0.068 ± $0$ $0$ 0.206 ± 0.008 ± $0$ 0.386 ± 0.026 ± 0.002 ± 0.336 ± 0.080 ± 0.002 ± 0.142 ± 0.058 ± 0.002
STORM-w/o-TS 0.943 0.250 0.014 1.587 0.194 0.011 1.223 0.174 0.010 0.203 0.418 0.018 0.800 0.336 0.015
± 0.354 ± 0.062 ± 0.001 ± 0.296 ± 0.038 ± 0.001 ± 0.236 ± 0.014 ± 0.001 ± 0.284 ± 0.095 ± 0.002 ± 0.063 ± 0.044 ± 0.002
STORM-w/o-CS 0.605 0.246 0.013 0.858 0.327 0.013 1.426 0.157 0.010 0.112 0.491 0.020 0.959 0.363 0.017
± 0.114 ± 0.115 ± 0.004 ± 0.236 ± 0.064 ± 0.001 ± 0.509 ± 0.046 ± 0.004 ± 0.231 ± 0.161 ± 0.005 ± 0.092 ± 0.033 ± 0.001
Improvement(%) -10.531 -2.929 -30.000 19.144 8.491 -10.000 1.153 $0$11.111 14.943 -118.792 -80.000 -13.816 -71.429 -36.364
Table 3. Algorithmic trading task results on all models across six standard metrics (averaged over 5 runs).
5.2. Implement Details
We implement all experiments by using PyTorch and conduct the experiments on two 80GB memory NVIDIA RTX H100 GPUs. The code is provided in the repository  [^4] https://github.com/DVampire/Storm. The input data is organized into batches with dimensions $(B,W,N,D)$, where $B=16$ denotes the batch size. $N$ represents the number of stocks, 28 for the DJ30 index and 408 for the SP500 index. The dimension $W=64$ corresponds to the number of historical data days, and $D=152$ represents the total number of features, including the OHLC data, technical indicators, and temporal information.
For the time-series module, we use an 8-day period for each stock as a patch, meaning the patch size is $(8,1,D)$. In contrast, for the cross-sectional module, we use $N$ stocks for each day as a patch, so the patch size is $(1,N,D)$. For both the encoders and decoders, we employ stacked transformer blocks. Each encoder consists of 4 layers of blocks with 4 attention heads, while each decoder consists of 2 layers of blocks with 8 attention heads. The size of both codebooks is selected from the set {256, 512, 1024}, with 512 yielding the best results. A consistent embedding dimension of 256 is employed across the encoders, decoders, and codebooks.
For the optimizer, we use AdamW with a learning rate of $1\times 10^{-4}$ and a weight decay of 0.05. The learning rate starts at 0, ramps up to $1\times 10^{-4}$ over 100 epochs with a linear warmup scheduler, and then decays to 0 over 1000 epochs. For the loss function balancing, the coefficients for the clip loss, as well as the price reconstruction losses for both the TS and CS modules, are all set to $1\times 10^{-3}$. The commitment loss coefficient for the codebooks is 1.0, while the coefficients for the orthogonality loss and diversity loss are both set to 0.1. Additionally, the coefficient for the price prediction loss of the factor module, as well as the KL divergence loss between the posterior and prior distributions, is also set to 0.1.
For downstream tasks, we employ the TopK-Drop rule strategy for portfolio management, using the recommended parameters from Qlib, where $k=5$ and $d=3$. For the trading task, we adopt the Proximal Policy Optimization (PPO) RL strategy. The CS and TS factors from the pretraining phase are used as the state. Notably, we retain all CS factors, which represent market characteristics, while only preserving the TS factors relevant to the currently traded stocks, reflecting stock-specific characteristics. The action space has a dimension of 3, corresponding to the choices of buy, hold, and sell. Our reward calculation considers value to account for scenarios where remaining cash may not cover the purchase of a new stock. Relying solely on price changes without considering position and cash factors is unreasonable. Value is defined as $v=p\times m+c$ where $p$ is the current price, $m$ is the position, and $c$ is the remaining cash. The reward is calculated as $r=\frac{v_{\text{post}}-v_{\text{pre}}}{v_{\text{pre}}}.$ Unlike compound interest, our method measures the rate of change from the initial value to the current value. For both downstream tasks, we start with an initial cash of $1\times 10^{6}$, while also accounting for a transaction cost ratio of $1\times 10^{-4}$.
For the pretraining prediction tasks and the downstream portfolio management tasks, each set of experiments is conducted 10 times with different random seeds, and the mean and variance are reported. For the trading tasks, each set of experiments is conducted 5 times with different random seeds, and the mean of the evaluation metrics is reported. Notably, the default parameters provided by the official implementations are used for baseline models, ensuring a consistent and fair comparison across different approaches.
5.3. Performance on Factor Quality (for RQ1)
To demonstrate the effectiveness and quality of the factors, we designed a stock future return prediction task to showcase the factors’ predictive ability and robustness to capture dynamic market changes. In the prediction task, STORM shows substantial improvements over six baseline methods across the SP500 and DJ30 datasets. Specifically, STORM achieves an average improvement of 16.105% on SP500 and 16.399% on DJ30, indicating the learned latent factors are significantly effective in capturing stock price trends.
The diversity of factors is crucial in financial investment, as it ensures the model captures the multi-dimensional characteristics of the market, avoiding redundancy or over-reliance on a single factor. In the codebook, each embedding represents a cluster center for a category of factor embeddings. To analyze the utilization rate of each factor category, we report the frequency of codebook embedding indices in the testing data. For clarity, we group every five indices into one interval, as shown in Figure 3(a). The results indicate that the distribution of factor usage is relatively even, demonstrating that the model has achieved good diversity in the factor selection process.
Figure 3. (a) Distribution of counts for codebook embedding indices. (b) the RankIC and RankICIR results across different codebook sizes.
We experiment with different codebook sizes on the DJ30 dataset to find the optimal balance between factor diversity and representation quality. If the codebook is too large, it may lead to sparse factor representation; if too small, it may result in insufficient factor differentiation and information redundancy. As shown in Figure 3(b), codebook size of 512 achieves optimal factor prediction performance.
5.4. Performance on downstream tasks (for RQ2)
We use three types of metrics: profit, risk-adjusted profit, and risk to evaluate STORM’s investment performance on downstream tasks, i.e., portfolio management and algorithmic trading. For profitability, STORM outperforms all the baselines, with average improvements of 106.10% on APY over the best-performing baseline in the PM task and 10.45% over the best baseline methods on three stocks in the AT task. For risk resistance, despite the test periods for both datasets, including the COVID-19 event, STORM still achieves the best AVO value, surpassing all comparison methods in the PM task. Moreover, considering the return and risk simultaneously, STORM performs the best on ASR with an average improvement of 58.64% over the best baseline result in the PM task, and surpasses the best baseline results across three stocks in the AT task. Such an outstanding ability to balance return and risk is significant in real investment practice.
5.5. Notations
We provide the main notations used in the paper.
Notation Description
$N$ total number of stocks
$W$ window size of the observation interval
$D$ total number of features of data
$T$ total number of trading days
$K$ codebook size
$p_{i,t}$ price of stock $i$ at day $t$
$\mathbf{x}$ observed data, which contains stock historical price data $\mathbf{p}$ and indicators $\mathbf{d}$
$\mathbf{x}^{\prime}_{ts}$, $\mathbf{x}^{\prime}_{cs}$, reconstructed observed data of the TS and CS modules
$\hat{\mathbf{y}}_{i,t}$, $\mathbf{y}_{i,t}$ predicted and true return of stock $i$ at day $t$
$\mathbf{w}_{t}$ the portfolio weight vector for $N$ stocks at trading day $t$
$a_{i,t}$ action of day $t$ on the stock $i$
$s_{t}$ state of day $t$
$\mathbf{z}_{e}^{ts}(\mathbf{x})$, $\mathbf{z}_{e}^{cs}(\mathbf{x})$ latent embeddings of the TS and CS modules
$\mathbf{z}_{q}^{ts}(\mathbf{x})$, $\mathbf{z}_{q}^{cs}(\mathbf{x})$ quantized embeddings of the TS and CS modules
$\mathbf{z}_{e}(\mathbf{x})$ factor embeddings
$\mathbf{z}$ posterior or latent factors
$k$ total number of selected stocks in PM strategy
$d$ the max number of stocks to be sold out
Table 4. Notations in the paper.
5.6. Ablation Study (for RQ3)
We conducted ablation experiments with two simplified versions of STORM, namely STORM-w/o-TS and STORM-w/o-CS, which only extract cross-sectional and time-series factors, respectively. Specifically, the absence of either time-series or cross-sectional factors significantly reduces the model’s effectiveness across various tasks. In the stock return prediction task, STORM outperforms both simplified models, with an average improvement of 21.73% on the RankIC metric and 24.54% on the RankICIR metric, highlighting the enhanced accuracy achieved by incorporating both time-series and cross-sectional factors. In the PM and AT tasks, STORM and two simplified models all performed well. Notably, STORM-w/o-TS exhibited the lowest AVO among the three models, although its profitability did not surpass that of STORM. These findings suggest that methods like FactorVAE (Duan et al., 2022), which focus solely on cross-sectional factors, may struggle to capture the full complexity of the financial market. Therefore, integrating both time-series and cross-sectional factors is essential for robust and accurate financial modeling.
6. Conclusions and Future Work
In this paper, we propose STORM, a spatio-temporal latent factor model that leverages the dual VQ-VAE framework to learn both time-series and cross-sectional factors. STORM ensures the orthogonality and diversity of factors, which are crucial for effective factor selection in financial trading. Extensive experiments demonstrate the superiority of STORM over the state-of-the-art methods across a prediction task and two downstream tasks, showcasing its robust performance and effectiveness. In the future, we plan to extend STORM by incorporating more side information, e.g., news, to explore the effects of exogenous factors.

### v3

@@
 The extensive experiments demonstrate STORM’s flexibility in adapting to downstream tasks and superior performance over baseline models.
-Code is available in PyTorch [^1] https://github.com/DVampire/Storm.
+Code is available in PyTorch [^1] https://anonymous.4open.science/r/Storm-2823.
 Factor Model, VQVAE, Portfolio Management, Algorithmic Trading
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the ACM Web Conference 2025; April 28–May 02, 2025; Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Computing methodologies Machine learningccs: Applied computing Electronic commerce
+conference: Proceedings of the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining 2025; August 03–August 07, 2025; Toronto, Canada
 1. Introduction
@@
 In order to address the challenges, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders (VQ-VAE), named STORM, with the architecture shown in Figure 1.
-We represent the factors as multi-dimensional embeddings to capture the complexity and nonlinearity of financial data, thereby improving the factors’ ability to account for factor returns (CH1).
+The model encodes factors as multi-dimensional embeddings, effectively capturing the complexity and nonlinearity inherent in financial data, and enhancing the capacity of factors to account for factor returns (CH1).
 Additionally, we develop a dual VQ-VAE architecture to capture cross-sectional and time-series features, considering both spatial and temporal perspectives [^2] We treat cross-sectional as spatial and time-series as temporal.
 These terms will be used interchangeably unless specified..
-We integrate cross-sectional and time-series features at both fine-grained and semantic levels to construct effective factors.
-To further enhance the model, we introduce diversity loss and orthogonality loss to ensure factor embedding diversity and independence (CH2).
-Furthermore, we treat the embeddings in the codebook as cluster centers for factor embeddings, effectively using them as class tokens to distinguish between factors.
-This approach makes the factor differentiation and selection process clear and transparent (CH3).
+By integrating these features at both fine-grained and semantic levels, the model constructs more effective factors.
+Through diversity and orthogonality loss constraints, we ensure representation independence among learned representations (CH2).
+Furthermore, codebook embeddings act as cluster centers, serving as class tokens to categorize factor embeddings.
+This strategy provides clarity and transparency to the differentiation and selection process of factors (CH3).
 Specifically, our contributions are three-fold:
-  • We design a dual VQ-VAE architecture to construct cross-sectional and time-series factors from both spatial and temporal perspectives.
+  • We design a dual VQ-VAE architecture that systematically constructs cross-sectional and time-series factors by capturing both spatial and temporal dependencies.
   • We leverage vector quantization techniques to improve factor embedding representation by reducing noise, enhancing diversity, and ensuring orthogonality.
 This approach significantly strengthens the factors’ predictive power and a more transparent factor selection process.
-  • Experiments on two U.S.
-stock markets across the stock future return prediction task and two downstream tasks (i.e., portfolio management and algorithmic trading) show that STORM outperforms all baselines across 2 prediction metrics and 6 standard financial metrics.
+  • We develop STORM as a versatile all-rounder capable of supporting three tasks simultaneously, whereas other methods tend to be specialized in specific tasks.
+  • Experiments on seven datasets from two U.S.
+stock markets, covering the stock future return prediction task (DJ30 and SP500), portfolio management (DJ30 and SP500), and algorithmic trading (five stocks), show that STORM outperforms all baselines across two prediction metrics and six standard financial metrics.
 2. Related Work
@@
 Recently, latent factor models have emerged that improve investment returns by studying latent factors self-adaptively.
-Specifically, (Gu et al., 2021) learns latent factors depend on assets characteristics, (Duan et al., 2022) learns cross-sectional factors by predicting future returns, and (Wei et al., 2023) learns market-augmented latent factors.
+Specifically, (Gu et al., 2021) learns latent factors depend on asset characteristics, (Duan et al., 2022) extracts cross-sectional factors by predicting future returns, and (Wei et al., 2023) learns market-augmented latent factors.
 However, these VAE-based latent factor models may suffer from a lack of robustness in representing low signal-to-noise ratio markets, leading to reduced reconstruction and generation capabilities.
-Additionally, their effectiveness is constrained by a primary focus on with insufficient attention to the temporal perspective.
+Furthermore, their effectiveness is constrained by an insufficient focus on the temporal perspective.
 2.2. Financial Trading
@@
 Quantitative traders leverage mathematical models and algorithms to automatically identify trading opportunities (Sun et al., 2024).
-Within the Fin-tech field, research tasks include portfolio management (PM), algorithmic trading (AT), order execution (Fang et al., 2021), market making (Spooner and Savani, 2020), etc.
+Within the FinTech domain, key research areas encompass portfolio management (PM), algorithmic trading (AT), order execution (Fang et al., 2021), market making (Spooner and Savani, 2020), etc.
+In this work, we focus on PM and AT, as they are fundamental to developing intelligent financial decision-making systems.
 PM focuses on optimizing wealth allocation across assets (Li and Hoi, 2014).
 Its development has evolved from simple rule-based methods (Poterba and Summers, 1988) to prediction-based approaches using machine learning (Ke et al., 2017) and deep learning (Qin et al., 2017; Zhao et al., 2022) for predicting returns or price movements.
-Despite recent integration of reinforcement learning (RL) in PM (Zhang et al., 2024b), the latent factor models still rely on deep learning for market modeling and precise predictions.
-AT involves trading financial assets using algorithm-generated signals.
-Similar to PM, there are rule-based and supervised learning methods (Yang et al., 2020) in the AT task.
-In recent years, RL-based methods (Sun et al., 2022) have become dominant for their superiority in complex sequential decision-making.
+Despite recent integration of reinforcement learning (RL) in PM (Zhang et al., 2024b), the latent factor models still predominantly rely on deep learning for market modeling and precise predictions.
+AT, on the other hand, involves executing trades with algorithm-generated signals.
+Like PM, AT has seen a transition from rule-based and supervised learning techniques (Yang et al., 2020) to RL-driven approaches.
+In particular, RL-based methods (Sun et al., 2022) have gained prominence due to their ability to navigate complex sequential decision-making problems, further enhancing trade execution efficiency and adaptability in dynamic market environments.
 2.3. Vector Quantized Variational Autoencoder
-VQ-VAE has made significant contributions across several areas of deep learning, particularly in representation learning, generative modeling, and time series analysis.
-Initially introduced by Van Den Oord et al., VQ-VAE has showcased its exceptional capabilities in processing high-dimensional data, e.g., images (Razavi et al., 2019), audio (Zhang et al., 2024a), and video (Yan et al., 2021), with a strong emphasis on learning discrete latent representations.
-Moreover, VQ-VAE has proven highly effective in handling complex time-dependent data (Talukder et al., 2024), advancing the development of robust generative models specifically in the context of temporal analysis.
+Vector Quantized Variational Autoencoder (VQ-VAE) has emerged as a fundamental framework in deep learning research, establishing methodological significance across core domains: representation learning, generative modeling, and time series analysis.
+Initially introduced by Van Den Oord et al., VQ-VAE has showcased its capabilities in processing high-dimensional data across diverse modalities, including images (Razavi et al., 2019), audio (Zhang et al., 2024a), and video (Yan et al., 2021), with a strong emphasis on learning discrete latent representations.
+Particularly noteworthy is the model’s demonstrated proficiency in capturing temporal dependencies, where recent studies like (Talukder et al., 2024) have validated its superior performance in modeling sequential patterns compared to conventional continuous latent space approaches, thereby advancing the development of robust generative architectures specifically optimized for temporal dynamics analysis.
 3. Preliminaries
@@
 (1) $\mathbb{E}[R_{i}^{e}]=\alpha_{i}+\beta_{i}\lambda\text{.}$
-$\mathbb{E}[\cdot]$ denotes the expectation operator.
-$\mathbb{E}[R_{i}^{e}]$ is the expected excess return.
+Here, $\mathbb{E}[\cdot]$ denotes the expectation operator, and $\mathbb{E}[R_{i}^{e}]$ represents the expected excess return of asset $i$.
 $\beta_{i}$ is factor exposure (or factor loading) of asset $i$, and $\lambda$ is factor expected return (or factor risk premium).
-$\alpha_{i}$ is pricing error.
-According to (Duan et al., 2022), they define the factor model in the spatial perspective, which only focuses on why the expected returns of different assets vary, rather than how the returns of each asset change over time.
+Lastly, $\alpha_{i}$ captures pricing error, reflecting any deviations from the model’s predicted returns.
+According to (Duan et al., 2022), the factor model is formulated from a spatial perspective, primarily examining cross-sectional variations in expected returns across different assets, rather than capturing the temporal evolution of individual asset returns.
 Therefore, we expand it from the temporal perspective and introduce the general functional form of the latent factor model:
@@
 From the temporal perspective, Equation 2 can be expressed as $\mathbf{y}_{t}=\bm{\alpha}_{t}+\sum_{k=1}^{K}\bm{\beta}_{t}^{k}\mathbf{z}_{t}^{k}+\bm{\epsilon}_{t}$, which we regard as the time-series factor model.
-The cross-sectional factor model focuses on the differences in returns of different assets at the same point in time, while the time-series factor model focuses on the changes in returns of a single asset at different points in time.
+The cross-sectional factor model focuses on the differences in returns of different assets at the same point in time, while the time-series factor model focuses on the changes in returns of a single asset at different time points.
 3.2. Problem Formulation
 To assess the effectiveness and generating ability of factors, we predict and evaluate the stock future returns.
-To show the adaptability of the latent factor model with multiple downstream tasks, we employ two financial trading tasks, i.e., portfolio management (Wang et al., 2021; Zhang et al., 2024b; Ye et al., 2020) and algorithmic trading (Sun et al., 2022; Liu et al., 2020), to showcase the investment capabilities.
-Observed Data: we utilize the stock’s historical price data $\mathbf{p}$(i.e., open, high, low, and close) and technical indicators $\mathbf{d}$ as observed variables $\mathbf{x}:=[\mathbf{p},\mathbf{d}]\in\mathbb{R}^{N\times W\times D}$, within a window size of $W$.
+To show the adaptability of the latent factor model with multiple downstream tasks, we employ two financial trading tasks, i.e., portfolio management (Wang et al., 2021a; Zhang et al., 2024b; Ye et al., 2020) and algorithmic trading (Sun et al., 2022; Liu et al., 2020), to showcase the model’s investment capabilities.
+Observed Data.
+We utilize the stock’s historical price data $\mathbf{p}$ (i.e., open, high, low, and close) and technical indicators $\mathbf{d}$ as observed variables $\mathbf{x}:=[\mathbf{p},\mathbf{d}]\in\mathbb{R}^{N\times W\times D}$, within a window size of $W$.
 The historical price data of stock $i$ at time $t$ is denoted as $\mathbf{p}_{i,t}:=[p_{i,t}^{o},p_{i,t}^{h},p_{i,t}^{l},p_{i,t}^{c}]\in\mathbb{R}^{D_{1}}$.
@@
 Each stock is represented by $D=D_{1}+D_{2}$ features per trading day, which we then use to predict the stock’s future returns, defined as $\mathbf{y}_{i,t+1}:=\frac{p_{i,t+1}-p_{i,t}}{p_{i,t}}$.
-Downstream tasks:
-  • Portfolio management task aims to construct a portfolio based on the predicted returns to test the profitability of the factor model.
+Downstream tasks. These primarily encompass the two most widely studied areas in quantitative finance – Portfolio management and Algorithmic trading:
+  • Portfolio management task aims to construct an optimal portfolio based on the predicted returns to test the profitability of the factor model.
 At each trading day $t$, we obtain the portfolio weight $\mathbf{w}_{t}=[w_{t}^{0},w_{t}^{1},...,w_{t}^{N-1}]\in\mathbb{R}^{N}$ for $N$ stocks.
-Each stock $i$ is assigned a weight $w_{t}^{i}$ representing its portfolio proportion, subject to the constraint that $\sum_{i=0}^{N-1}w_{t}^{i}=1$ for full investment.
+Each stock $i$ is assigned a weight $w_{t}^{i}$ representing its portfolio proportion within the portfolio.
+To ensure full capital deployment, the allocation must satisfy the budge constraint $\sum_{i=0}^{N-1}w_{t}^{i}=1$.
   • Algorithmic trading task aims to execute buy, hold, and sell actions based on predicted asset states to balance returns and risks.
@@
 4.1. Dual VQ-VAE Structure
-The dual VQ-VAE architecture processes stock data from spatial and temporal perspectives by patching the data in both modules.
-Then, the encoder extracts time-series (TS) and cross-sectional (CS) features, which are subsequently clustered using a learnable codebook for quantized representation.
-Finally, the decoder reconstructs the data, capturing complex patterns and enhancing forecasting performance.
+The dual VQ-VAE architecture processes stock data from both spatial and temporal perspectives by patching the data within each module.
+The encoder then extracts time-series (TS) and cross-sectional (CS) features, which are subsequently clustered using a learnable codebook for quantized representation.
+Finally, the decoder reconstructs the data, effectively capturing complex patterns and enhancing model’s forecasting performance.
 4.1.1. Patching
@@
 Considering the Transformer’s (Vaswani et al., 2017) remarkable ability to capture both local and global dependencies via self-attention and patching strategies, it is particularly well-suited for long-term time-series forecasting (Nie et al., 2022).
-Therefore, we employ stacked transformer blocks as the encoders in the TS and CS modules to capture complex time-series and cross-sectional patterns, ultimately improving the accuracy and robustness of feature extraction.
+Therefore, we employ stacked transformer blocks as the encoders in both the TS and CS modules to capture complex time-series and cross-sectional patterns.
+This architecture enhances the accuracy and robustness of feature extraction, ultimately improving forecasting performance.
+While transformers are widely applicable across domains such as natural language processing (NLP) and computer vision (CV), STORM’s flexible encoder-decoder framework allows for the integration of alternative architectures, such as long short-term memory (LSTM) (Wang et al., 2021b) and temporal convolutional networks (TCN) (Yu and Koltun, 2015).
 4.1.3. Codebook Construction and Optimization
@@
 ii) Explicit Factor Selection: The discrete token indices facilitate an explicit factor selection process, identifying the most relevant factors that influence stock returns and thus enhancing prediction accuracy.
-iii) Noise Reduction: The factor selection process filters out irrelevant or redundant information, reducing noise and improving the overall robustness of the model.
+iii) Noise Reduction: By filtering out irrelevant or redundant information during factor selection, VQ-VAE mitigates noise, contributing to improved model robustness and more reliable financial insights.
 To encourage equal usage of the codebook vectors (Baevski et al., 2020), we incorporate a diversity loss to ensure balanced representations across the codebook.
 It is achieved by maximizing the entropy of the averaged softmax distribution over vectors for each codebook $\bar{p}_{g}$:
-(4) $\mathcal{L}_{\text{div}}=\frac{1}{GK}\sum_{g=1}^{G}\sum_{k=1}^{K}\bar{p}_{g,k}\log{\bar{p}_{g,k}}\text{,}\vspace{-0.2cm}$
+(4) $\mathcal{L}_{\text{div}}=\frac{1}{GK}\sum_{g=1}^{G}\sum_{k=1}^{K}\bar{p}_{g,k}\log{\bar{p}_{g,k}}\text{,}\vspace{-0.1cm}$
 where $G$ represents the number of codebooks, set 2 in STORM. The diversity constraint helps enhance representational capacity, leading to better coverage of the input space.
-However, the lack of independence among factors can lead to multicollinearity, causing the model to overfit and are less robust to various market situations.
+However, the lack of independence among factors can introduce multicollinearity, making the model prone to overfitting and reducing its robustness across different market conditions.
 To address this issue, we impose an orthogonality constraint on the codebook, as recent studies (Shin et al., 2023) have shown that enforcing orthogonality allows discretized codes to maintain translation equivariance.
@@
 We utilize Transformer as decoders either, to generate constructed data $\mathbf{x}_{ts}^{\prime}$ and $\mathbf{x}_{cs}^{\prime}$, which aim to closely approximate the original input data.
-Finally, the encoders, decoders, and codebooks are jointly learned by minimizing the following loss objectives:
+Finally, the encoders, decoders, and codebooks are jointly optimized by minimizing the following loss objectives:
 (6) $\mathcal{L}_{1}=$ $\lambda_{\text{ortho}}\mathcal{L}_{\text{ortho}}+\lambda_{\text{div}}\mathcal{L}_{\text{div}}+\left|\left|\mathbf{x}-\mathbf{x}^{\prime}_{ts}\right|\right|^{2}_{2}+\left|\left|\mathbf{x}-\mathbf{x}^{\prime}_{cs}\right|\right|^{2}_{2}$
@@
 The factor module aims to integrate the TS and CS latent features through feature fusion and alignment.
-Then, through prior-posterior learning to generate factor embeddings from latent features to predict stock future returns.
+Afterward, it utilizes prior-posterior learning to generate factor embeddings from latent features to predict stock future returns.
 4.3.1. Feature Fusion and Alignment
@@
 However, these methods result in two parallel processes that do not align at the feature level.
-To address this limitation, we employ a cross-attention mechanism to enhance the interaction and fusion of features at a fine-grained level, improving the model’s dynamic understanding of the input data.
+To address the limitation, we employ a cross-attention mechanism to enhance the interaction and fusion of features at a fine-grained level, improving the model’s dynamic understanding of the input data.
 Additionally, contrastive learning is used at the semantic level to enhance feature representation by emphasizing semantic similarities and differences.
@@
 In order to fuse the features in the TS and CS modules at a fine-grained level, we leverage a multi-scale encoder to combine information between two patches and produce stronger fusion features for factor generations, motivated by CrossViT (Chen et al., 2021).
-The encoder contains cross-attention layers that achieve both accuracy and efficiency.
+The architecture incorporates cross-attention layers, balancing both accuracy and computational efficiency, ensuring that factors derived from the combined features are more representative of market dynamics.
 Contrastive Learning.
-We incorporate contrastive learning, inspired by its success in aligning text and image features through a shared embedding space (Li et al., 2023; Radford et al., 2021), to enhance semantic consistency between time-series and cross-sectional features.
-In the factor module, contrastive loss ensures that similar features cluster together while dissimilar features are pushed apart in the embedding space.
+To further enhance semantic consistency between TS and CS features, we integrate contrastive learning, drawing inspiration from its success in aligning multimodal representations (Li et al., 2023; Radford et al., 2021).
+Within the factor module, a contrastive loss encourages similar features to cluster closely while pushing apart dissimilar ones in the embedding space.
 4.3.2. Prior-Posterior Learning
 Through multi-scale encoder and contrastive learning layer, the CS and TS latent features are fused and aligned.
-In order to retain the codebook’s categorical properties for factors, we add the codebook embeddings $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$ as extra [CLS]-tokens (Dosovitskiy et al., 2020) to the latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$, respectively.
+In order to retain the codebook’s categorical properties for factors, we incorporate the codebook embeddings $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$ as extra [CLS]-tokens (Dosovitskiy et al., 2020), which are appended to the latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$, respectively.
 Then, two latent features concat together to get the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$.
-Due to stock market’s inherent volatility and complexit, it is difficult to bridge the gap between the noisy market and effective factor model (Duan et al., 2022).
-Therefore, we utilize the prior-posterior structure to predict future returns and optimize latent factor embeddings.
-In the training stage, the posterior layer predict posterior distribution of factor expected returns $\mathbf{z}_{post}$ from the true future stock returns $\mathbf{y}$ and the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$:
+Given the inherent volatility and complexity of the stock market, bridging the gap between noisy market data and an effective factor model remains a significant challenge (Duan et al., 2022).
+Therefore, we introduce the prior-posterior structure to predict future returns and optimize latent factor embeddings.
+In the training stage, the posterior layer estimates posterior distribution of factor expected returns $\mathbf{z}_{post}$ from the true future stock returns $\mathbf{y}$ and the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$:
 (7) $[\mu_{post},\sigma_{post}]=\phi_{FE}(\mathbf{y},\mathbf{z}_{e}(\mathbf{x}))\text{,}$
@@
 (8) $[\mu_{prior},\sigma_{prior}]=\phi_{FP}(\mathbf{z}_{e}(\mathbf{x}))\text{.}$
-The return predictor uses factor expected return $\mathbf{z}$ to calculate future returns:
+The return predictor computes future returns based on the factor expected return $\mathbf{z}$:
 (9) $\hat{\mathbf{y}}=\bm{\alpha}+\sum_{k=1}^{K}\bm{\beta}^{k}\mathbf{z}^{k}+\bm{\epsilon}\text{.}$
+The predicted returns $\hat{\mathbf{y}}$ can be directly integrated into portfolio optimization, algorithmic trading strategies, enabling dynamic rebalancing and improved risk management.
 4.4. Downstream Tasks
@@
 RQ3: How do the key components contribute to the performance of STORM?
+5.1. Experiment Settings
+Dataset.
+We conduct experiments on two U.S.
+stock markets, SP500 and DJ30, using stock daily data that includes technical features based on Alpha158 (Yang et al., 2020).
+Both datasets span 16 years, from 2008-04-01 to 2024-03-31, encompassing global conditions, e.g., the 2007-2008 financial crisis and COVID-19.
+Datasets are chronologically divided into non-overlapping training (from 2008-04-01 to 2021-03-31) and test (from 2021-04-01 to 2024-03-31) sets.
+Metrics.
+We compare STORM and baselines in terms of 6 financial metrics across the PM and AT tasks.
+The financial metrics include 2 profit criteria, Annualized Percentage Yield (APY) and Cumulative Wealth (CW), 2 risk-adjusted profit criteria, Calmar Ratio (CR) and Annualized Sharpe Ratio (ASR), and 2 risk criteria, Maximum Drawdown (MDD) and Annualized Volatility (AVO).
+The calculation formulas and meanings of these metrics are as follows:
+  • CW is the total returns yielded from a portfolio strategy: $\text{CW}_{T}=\prod_{i=1}^{T}(1+r_{i})$, where $r_{i}$ is the net return.
+  • APY measures the average wealth increment that one portfolio strategy could achieve compounded in a year, which is defined as $\text{APY}_{T}=\sqrt[y]{\text{CW}_{T}}-1$, where $y$ is the number of years corresponding to $T$ trading rounds.
+  • MDD measures the largest loss from a historical peak in the cumulative wealth to show the worst case, which is defined as: $MDD=\mathop{\max}_{i=0}^{T}\frac{P_{i}-R_{i}}{P_{i}}$, where $R_{i}=\prod_{i=1}^{T}{\frac{V_{i}}{V_{i-1}}}$ and $P_{i}=\mathop{\max}_{i=1}^{T}R_{i}$.
+  • AVO is the annualized standard deviation of daily returns and multiplied by $\sqrt{AT}$, where $AT$ is the average trading rounds of annual trading days and $AT=252$ for all the datasets.
+  • ASR is an annualized volatility risk-adjusted return defined as $\text{ASR}=\frac{\text{APY}-R_{f}}{\text{AVO}}$, where $R_{f}$ is the risk-free return.
+  • CR measures the drawdown risk-adjusted return of a portfolio calculated as $CR=\frac{\text{APY}}{\text{MDD}}$.
+Typically, to evaluate the effectiveness of the learned factors in the PM task, we adopt the Rank Information Coefficient (RankIC) and the Information Ratio of RankIC (RankICIR):
+  • RankIC is a ranking metric in finance, which measures the correlation between the predicted rankings and the actual returns. It is defined as:
+    (10) $\text{RankIC}_{s}=\frac{1}{N}\frac{(r_{\hat{y}_{s}}-\text{mean}(r_{\hat{y}_{s}}))^{T}(r_{y_{t}}-\text{mean}(r_{y_{s}}))}{\text{std}(r_{\hat{y}_{s}})\cdot\text{std}(r_{y_{s}})}$
+    $\text{RankIC}=\frac{1}{T_{\text{test}}}\sum_{s=1}^{T_{\text{test}}}\text{RankIC}_{s}\text{,}$
+    where $T_{test}$ denotes the number of trading days in the test range, $r_{y_{s}}$ and $r_{\hat{y}_{s}}$ represent the true and predicted ranks of stocks on the trading day $s$, respectively.
+  • RankICIR is the information ratio of RankIC, which measures the stability of prediction, $\text{RankICIR}=\frac{\text{mean}(\text{RankIC}_{s})}{\text{std}(\text{RankIC}_{s})}$.
+Baselines.
+As shown in Table 2, we compare STORM with nine methods across one prediction task and two downstream tasks.
+The baseline methods can be divided into three categories: ML & DL-based models, Factor models, and RL-based models.
+We evaluate STORM against ML&DL models and factor models in prediction and PM tasks, and against ML&DL models and RL methods in the AT task.
+Moreover, STORM is a versatile all-rounder that supports three tasks, unlike other methods specialized in specific tasks.
+The following will provide a brief introduction to each method:
+  • Market
+    • Buy-and-Hold (B&H) involves holding assets for an extended period, regardless of short-term market fluctuations, assuming that long-term returns will be more favorable.
+  • ML&DL-based
+    • LGBM (Yang et al., 2020) uses a series of tree models to predict price fluctuations and provide buy and sell signals.
+    • LSTM (Yu and Li, 2018) utilizes long short-term memory to improve the accuracy of price predictions.
+    • Transformer (Yang et al., 2020) models leverage self-attention mechanisms to enhance the precision of price forecasts.
+  • RL-based
+    • SAC (Haarnoja et al., 2018) is an off-policy actor-critic algorithm that optimizes trading strategies using entropy regularization and soft value functions in continuous action spaces.
+    • PPO (Schulman et al., 2017) updates trading policies iteratively to balance exploration and exploitation, ensuring stability and sample efficiency.
+    • DQN (Mnih et al., 2013) uses deep neural networks to approximate the action-value function and make trading decisions from market data.
+  • Factor-model
+    • CAFactor (Gu et al., 2021) is a state-of-the-art model that learns latent factors depending on asset characteristics.
+    • FactorVAE (Duan et al., 2022) is a state-of-the-art VAE-based latent factor model that learns optimal latent stock factors.
+    • HireVAE (Wei et al., 2023) is a state-of-the-art VAE-based latent factor model that identifies factors between market situation and stock-wise.
 Models SP500 DJ30
@@
 Table 1. Results of all models on two metrics, i.e., RankIC and RankICIR (mean ± range, computed across 10 runs).
-5.1. Experiment Settings
-Dataset.
-We conduct experiments on two U.S.
-stock markets, SP500 and DJ30, using stock daily data that includes technical features based on Alpha158 (Yang et al., 2020).
-Both datasets span 16 years, from 2008-04-01 to 2024-03-31, encompassing global conditions, e.g., the 2007-2008 financial crisis and COVID-19.
-Datasets are chronologically divided into non-overlapping training (from 2008-04-01 to 2021-03-31) and test (from 2021-04-01 to 2024-03-31) sets.
-Metrics.
-We compare STORM and baselines in terms of 6 financial metrics across the PM and AT tasks.
-The financial metrics include 2 profit criteria, Annualized Percentage Yield (APY) and Cumulative Wealth (CW), 2 risk-adjusted profit criteria, Calmar Ratio (CR) and Annualized Sharpe Ratio (ASR), and 2 risk criteria, Maximum Drawdown (MDD) and Annualized Volatility (AVO).
-The calculation formulas and meanings of these metrics are as follows:
-  • CW is the total returns yielded from a portfolio strategy: $\text{CW}_{T}=\prod_{i=1}^{T}(1+r_{i})$, where $r_{i}$ is the net return.
-  • APY measures the average wealth increment that one portfolio strategy could achieve compounded in a year, which is defined as $\text{APY}_{T}=\sqrt[y]{\text{CW}_{T}}-1$, where $y$ is the number of years corresponding to $T$ trading rounds.
-  • MDD measures the largest loss from a historical peak in the cumulative wealth to show the worst case, which is defined as: $MDD=\mathop{\max}_{i=0}^{T}\frac{P_{i}-R_{i}}{P_{i}}$, where $R_{i}=\prod_{i=1}^{T}{\frac{V_{i}}{V_{i-1}}}$ and $P_{i}=\mathop{\max}_{i=1}^{T}R_{i}$.
-  • AVO is the annualized standard deviation of daily returns and multiplied by $\sqrt{AT}$, where $AT$ is the average trading rounds of annual trading days and $AT=252$ for all the datasets.
-  • ASR is an annualized volatility risk-adjusted return defined as $\text{ASR}=\frac{\text{APY}-R_{f}}{\text{AVO}}$, where $R_{f}$ is the risk-free return.
-  • CR measures the drawdown risk-adjusted return of a portfolio calculated as $CR=\frac{\text{APY}}{\text{MDD}}$.
-Typically, to evaluate the effectiveness of the learned factors in the PM task, we adopt the Rank Information Coefficient (RankIC) and the Information Ratio of RankIC (RankICIR):
-  • RankIC is a ranking metric in finance, which measures the correlation between the predicted rankings and the actual returns. It is defined as:
-    (10) $\text{RankIC}_{s}=\frac{1}{N}\frac{(r_{\hat{y}_{s}}-\text{mean}(r_{\hat{y}_{s}}))^{T}(r_{y_{t}}-\text{mean}(r_{y_{s}}))}{\text{std}(r_{\hat{y}_{s}})\cdot\text{std}(r_{y_{s}})}$
-    $\text{RankIC}=\frac{1}{T_{\text{test}}}\sum_{s=1}^{T_{\text{test}}}\text{RankIC}_{s}\text{,}$
-    where $T_{test}$ is the number of trading days of the test range, $r_{y_{s}}$ and $r_{\hat{y}_{s}}$ are the true and predicted ranks of stocks on the trading day $s$.
-  • RankICIR is the information ratio of RankIC, which measures the stability of prediction, $\text{RankICIR}=\frac{\text{mean}(\text{RankIC}_{s})}{\text{std}(\text{RankIC}_{s})}$.
-Baselines.
-We compare STORM with nine methods across one prediction task and two downstream tasks.
-The baseline methods can be divided into three categories: ML & DL-based models, Factor models, and RL-based models.
-In the prediction task and PM task, we compare STORM with ML & DL-based and factor models.
-In the AT task, we compare STORM with ML & DL-based and RL-based methods.
-We use the labels P, PM, and AT to indicate the applicable task scope for the following baselines.
-The following will provide a brief introduction to each method:
-  • Market
-    • Buy-and-Hold (B&H) involves holding assets for an extended period, regardless of short-term market fluctuations, assuming that long-term returns will be more favorable. (P, PM, AT)
-  • ML&DL-based
-    • LGBM (Yang et al., 2020) uses a series of tree models to predict price fluctuations and provide buy and sell signals. (P, PM)
-    • LSTM (Yu and Li, 2018) utilizes long short-term memory to improve the accuracy of price predictions. (P, PM)
-    • Transformer (Yang et al., 2020) models leverage self-attention mechanisms to enhance the precision of price forecasts. (P, PM)
-  • RL-based
-    • SAC (Haarnoja et al., 2018) is an off-policy actor-critic algorithm that optimizes trading strategies using entropy regularization and soft value functions in continuous action spaces. (PM, AT)
-    • PPO (Schulman et al., 2017) updates trading policies iteratively to balance exploration and exploitation, ensuring stability and sample efficiency. (PM, AT)
-    • DQN (Mnih et al., 2013) uses deep neural networks to approximate the action-value function and make trading decisions from market data. (PM, AT)
-  • Factor-model
-    • CAFactor is a state-of-the-art model that learns latent factors depending on asset characteristics. (P, PM)
-    • FactorVAE (Duan et al., 2022) is a state-of-the-art VAE-based latent factor model that learns optimal latent stock factors. (P, PM)
-    • HireVAE (Wei et al., 2023) is a state-of-the-art VAE-based latent factor model that identifies factors between market situation and stock-wise. (P, PM)
+Task Market ML & DL-based RL-based Factor model Ours
+B&H LGBM (Yang et al., 2020) LSTM (Yu and Li, 2018) Transformer (Yang et al., 2020) SAC (Haarnoja et al., 2018) PPO (Schulman et al., 2017) DQN (Mnih et al., 2013) CAFactor (Gu et al., 2021) FactorVAE (Duan et al., 2022) HireVAE (Wei et al., 2023) STORM
+Prediction
+Portfolio Management
+Algorithmic Trading
+  • * indicates that the model is applicable, while signifies that it is NOT applicable to the corresponding task.
+Table 2. Comparison of baselines across different tasks
+Implementation Details.
+All experiments are implemented using PyTorch and conducted on two NVIDIA RTX H100 GPUs with 80GB of memory each.
+The implementations of the ML & DL methods are based on Qlib (Yang et al., 2020).
+The input data is organized into batches with dimensions corresponding to historical days, stocks, and features, ensuring compatibility across tasks.
+As for other baselines, we adopt the default settings in their public implementations.
+Each experiment is repeated multiple times with different random seeds, and the mean and variance of the metrics are reported.
+Detailed information regarding the network architecture of STORM, as well as the optimizers and hyperparameters used for the two downstream tasks, is provided in Appendix B.
 SP500 Dataset
@@
 ± 0.073 ± 0.209 ± 0.213 ± 0.026 ± 0.071 ± 0.059 ± 0.047 ± 0.136 ± 0.205 ± 0.177 ± 0.514 ± 0.418 ± 0.081 ± 0.232 ± 0.105
-STORM 0.137 1.472 0.501 0.134 1.459 0.363 0.239 1.903 1.263 0.099 1.339 0.433 0.250 1.954 0.724
-± 0.015 ± 0.058 ± 0.035 ± 0.039 ± 0.151 ± 0.072 ± 0.033 ± 0.150 ± 0.224 ± 0.085 ± 0.321 ± 0.289 ± 0.034 ± 0.163 ± 0.086
-STORM-w/o-TS 0.124 1.428 0.714 0.127 1.437 0.627 0.126 1.428 0.703 -0.001 1.000 0.099 0.103 1.341 0.360
-± 0.062 ± 0.229 ± 0.003 ± 0.058 ± 0.214 ± 0.187 0± 0.022 ± 0.085 ± 0.091 ± 0.049 ± 0.152 ± 0.171 ± 0.017 ± 0.062 ± 0.065
-STORM-w/o-CS 0.059 1.187 0.309 0.103 1.347 0.350 0.149 1.525 0.908 -0.031 0.913 0.048 0.148 1.513 0.445
-± 0.021 ± 0.070 ± 0.061 ± 0.043 ± 0.154 ± 0.106 ± 0.095 ± 0.400 ± 0.450 ± 0.053 ± 0.156 ± 0.109 ± 0.03 ± 0.117 ± 0.040
-Improvement(%) -6.803 -2.452 35.227 0.752 5.419 2.117 5.286 15.054 7.949 25.316 3.798 -5.459 -6.367 -5.467 -22.067
+STORM 0.229 1.857 0.750 0.174 1.621 0.559 0.236 1.893 1.470 0.173 1.625 0.773 0.290 2.154 1.216
+± 0.033 ± 0.154 ± 0.066 ± 0.032 ± 0.133 ± 0.081 ± 0.039 ± 0.184 ± 0.445 ± 0.067 ± 0.284 ± 0.293 ± 0.052 ± 0.262 ± 0.597
+STORM-w/o-TS 0.199 1.730 0.766 0.127 1.437 0.627 0.154 1.536 1.294 0.106 1.212 0.267 0.254 1.979 0.964
+± 0.059 ± 0.260 ± 0.274 ± 0.058 ± 0.214 ± 0.187 ± 0.023 ± 0.089 ± 1.012 ± 0.120 ± 0.145 ± 0.135 ± 0.060 ± 0.093 ± 0.503
+STORM-w/o-CS 0.148 1.521 0.641 0.103 1.347 0.350 0.152 1.534 0.814 0.136 1.474 0.541 0.181 1.650 0.648
+± 0.069 ± 0.283 ± 0.184 ± 0.043 ± 0.154 ± 0.106 ± 0.054 ± 0.219 ± 0.352 ± 0.082 ± 0.300 ± 0.328 ± 0.051 ± 0.221 ± 0.361
+Improvement(%) 55.782 23.062 45.076 30.827 17.124 2.117 3.965 14.4501 20.408 118.987 26.969 66.594 8.614 4.159 30.893
 Models AAPL JPM IBM INTC MSFT
@@
 ± 0.472 ± 0.159 ± 0.007 ± 1.270 ± 0.118 ± 0.003 ± 0.403 ± 0.049 ± 0.004 ± 0.584 ± 0.310 ± 0.009 ± 0.148 ± 0.079 ± 0.004
-SAC 1.054 0.292 0.016 1.044 0.350 0.014 1.794 0.157 0.011 0.527 0.359 0.017 1.672 0.229 0.013
+SAC 1.054 0.292 0.016 1.044 0.350 0.014 1.794 0.157 0.011 0.527 0.359 0.017 1.638 0.229 0.013
 ± 0.223 ± 0.022 ± 0.002 ± 0.276 ± 0.085 ± 0.001 ± 0.458 ± 0.068 ± 0.005 ± 1.061 ± 0.191 ± 0.004 ± 0.736 ± 0.215 ± 0.005
@@
 ± 0.412 ± 0.030 ± 0.001 ± 0.189 ± 0.018 ± 0.001 ± 0.231 ± 0.027 ± 0.003 ± 1.689 ± 0.115 ± 0.007 ± 0.299 ± 0.036 ± 0.001
-STORM 0.913 0.309 0.017 0.983 0.389 0.015 1.843 0.162 0.011 0.700 0.326 0.019 1.441 0.312 0.016
-± 0.068 ± $0$ $0$ 0.206 ± 0.008 ± $0$ 0.386 ± 0.026 ± 0.002 ± 0.336 ± 0.080 ± 0.002 ± 0.142 ± 0.058 ± 0.002
-STORM-w/o-TS 0.943 0.250 0.014 1.587 0.194 0.011 1.223 0.174 0.010 0.203 0.418 0.018 0.800 0.336 0.015
-± 0.354 ± 0.062 ± 0.001 ± 0.296 ± 0.038 ± 0.001 ± 0.236 ± 0.014 ± 0.001 ± 0.284 ± 0.095 ± 0.002 ± 0.063 ± 0.044 ± 0.002
-STORM-w/o-CS 0.605 0.246 0.013 0.858 0.327 0.013 1.426 0.157 0.010 0.112 0.491 0.020 0.959 0.363 0.017
-± 0.114 ± 0.115 ± 0.004 ± 0.236 ± 0.064 ± 0.001 ± 0.509 ± 0.046 ± 0.004 ± 0.231 ± 0.161 ± 0.005 ± 0.092 ± 0.033 ± 0.001
-Improvement(%) -10.531 -2.929 -30.000 19.144 8.491 -10.000 1.153 $0$11.111 14.943 -118.792 -80.000 -13.816 -71.429 -36.364
+STORM 1.346 0.282 0.016 1.245 0.299 0.014 2.107 0.137 0.009 1.189 0.227 0.015 1.672 0.210 0.015
+± 0.162 ± $0$ 0.001 ± 0.106 ± 0.016 ± 0.001 ± 0.302 ± 0.033 ± 0.001 ± 0.386 ± 0.055 ± 0.002 ± 0.201 ± 0.067 ± 0.002
+STORM-w/o-TS 1.199 0.262 0.017 1.587 0.194 0.011 1.868 0.131 0.008 0.580 0.344 0.017 1.553 0.244 0.015
+± 0.300 ± 0.052 ± 0.001 ± 0.296 ± 0.038 ± 0.001 ± 0.956 ± 0.077 ± 0.003 ± 0.229 ± 0.061 ± $0$ 0.187 ± 0.093 ± 0.004
+STORM-w/o-CS 1.027 0.229 0.015 0.858 0.327 0.013 1.397 0.177 0.010 0.914 0.312 0.018 1.401 0.295 0.016
+± 0.326 ± 0.030 ± 0.001 ± 0.236 ± 0.064 ± 0.001 ± 0.405 ± 0.028 ± 0.001 ± 0.397 ± 0.136 ± 0.004 ± 0.224 ± 0.090 ± 0.002
+Improvement(%) 27.704 4.184 - 19.144 8.491 - 15.642 16.561 11.111 95.238 - - 2.076 - -
 Table 4. Algorithmic trading task results on all models across six standard metrics (averaged over 5 runs).
-5.2. Implement Details
+5.2. Factor Quality (for RQ1)
+To demonstrate the effectiveness and quality of the factors, we designed a stock future return prediction task to showcase the factors’ predictive ability and robustness to capture dynamic market changes.
+In the prediction task, STORM shows substantial improvements over six baseline methods across the SP500 and DJ30 datasets.
+Specifically, in terms of RankIC and RankICIR, STORM achieves an average improvement of 16.105% on SP500 and 16.399% on DJ30, as shown in Table  1.
+These results highlight the learned latent factors are significantly effective in capturing stock price trends.
+The diversity of factors is essential in financial investment, as it allows the model to capture the multi-dimensional characteristics of the market, avoiding redundancy or over-reliance on a single factor.
+In the codebook, each embedding represents a cluster center for a category of factor embeddings.
+To analyze the utilization rate of each factor category, we report the frequency of codebook embedding indices in the testing data.
+For clarity, we group every five indices into one interval, as illustrated in Figure 3(a).
+The results indicates a relatively even distribution of factor usage, highlighting the model’s effectiveness to maintain diversity during the factor selection process.
+Furthermore, we experiment with different codebook sizes on the DJ30 dataset to find the optimal balance between factor diversity and representation quality.
+A larger codebook may lead to sparse factor representations, reducing the effectiveness of factor utilization, while a smaller codebook could result in insufficient factor differentiation and information redundancy.
+Following standard practices, we evaluate five different codebook sizes, including 128, 256, 512, 1024, and 2048.
+As shown in Figure 3(b), codebook size of 512 achieves optimal factor prediction performance, striking a balance between expressiveness and efficiency.
+Figure 3. (a) Distribution of counts for codebook embedding indices. (b) the RankIC and RankICIR results across different codebook sizes.
+5.3. Downstream Task Performance (for RQ2)
+We use three types of metrics: profit, risk, and risk-adjusted profit to evaluate STORM’s investment performance on downstream tasks, i.e., portfolio management and algorithmic trading.
+For profitability, STORM outperforms all the baselines, with average improvements of 106.10% on APY over the best-performing baseline in the PM task and 43.64% over the best baseline methods on five stocks in the AT task.
+These results underscore the model’s capability to extract meaningful factors, and optimize decision-making for enhanced financial performance.
+For risk resistance, although STORM does not outperform the best baseline methods in terms of MDD and AVO on the INTC and MSFT datasets for the trading task, this can be attributed to the extreme price fluctuations during the COVID-19 event.
+Both stocks experienced a sharp decline of over 30%, followed by a subsequent rebound.
+Given STORM’s strong profitability capability, its trading strategy tends to aggressively capture trends, which naturally leads to larger drawdowns and increased volatility as a trade-off for higher potential returns.
+Moreover, considering the return and risk simultaneously, STORM performs the best on ASR and CR with an average improvement of 58.64% and 166.83% respectively, over the best baseline result in the PM task, and surpasses the best baseline results across three stocks in the AT task.
+Such an outstanding ability to balance return and risk is particularly valuable in real-world investment, where achieving high returns while effectively managing risk is crucial for long-term financial success.
+5.4. Ablation Study (for RQ3)
+We conducted ablation experiments with two simplified versions of STORM, namely STORM-w/o-TS and STORM-w/o-CS, which only extract cross-sectional and time-series factors, respectively.
+Specifically, the absence of either time-series or cross-sectional factors significantly reduces the model’s effectiveness across various tasks.
+In the stock return prediction task, STORM outperforms both simplified models, with an average improvement of 21.73% on the RankIC metric and 24.54% on the RankICIR metric, highlighting the enhanced accuracy achieved by incorporating both time-series and cross-sectional factors.
+In the PM and AT tasks, STORM and two simplified models all performed well.
+Notably, STORM-w/o-TS exhibited the lowest AVO among the three models, although its profitability did not surpass that of STORM.
+These findings suggest that methods like FactorVAE (Duan et al., 2022), which focus solely on cross-sectional factors, may struggle to capture the full complexity of the financial market.
+Therefore, integrating both time-series and cross-sectional factors is essential for robust and accurate financial modeling.
+5.5. Discussion on Performance
+Efficiency.
+STORM achieves remarkable efficiency in real-world trading scenarios.
+Unlike high-frequency trading systems requiring ultra-low latency and specialized hardware, our daily trading strategy operates efficiently on standard CPU hardware.
+This is enabled by a compact model architecture characterized by: i) a minimalist design with only a few transformer blocks (see Appendix B), and ii) a lightweight parameterization comprising 15.2 million FP32 parameters on DJ30.
+This design ensures seamless deployment on edge devices while supporting concurrent market analysis and trading.
+Effectiveness.
+STORM exhibits exceptional performance compared to baseline methods, showcasing versatility across tasks including PM and AT.
+Unlike baselines that often excel in specific tasks but lack generalizability, STORM consistently delivers robust overall results.
+Moreover, STORM can balance returns and risks by emphasizing comprehensive investment performance over individual financial metrics.
+Collectively, these attributes substantiate that the additional computational cost incurred by STORM is well justified by its robustness and adaptability across diverse tasks and datasets.
+Generalization.
+STORM demonstrates robust generalizability by achieving strong performance across markets with varying scales and distinct characteristics.
+It excels on both the DJ30, which reflects large-cap industrial blue chips, and the SP500, representing a broader market across diverse industries.
+Furthermore, its flexible encoder-decoder architecture allows the integration of additional layers such as Mamba  (Gu and Dao, 2023; Dao and Gu, 2024) and ensures applicability to diverse trading environments and asset classes.
+6. Conclusions and Future Work
+In this paper, we propose STORM, a spatio-temporal latent factor model that leverages the dual VQ-VAE framework to learn both time-series and cross-sectional factors.
+STORM ensures the orthogonality and diversity of factors, which are crucial for effective factor selection in financial trading.
+Extensive experiments demonstrate the superiority of STORM over the state-of-the-art methods across a prediction task and two downstream tasks, showcasing its robust performance and effectiveness.
+In the future, we plan to extend STORM by incorporating more side information, e.g., news, to explore the effects of exogenous factors.
+Appendix A Notations
+We provide the main notations in Table A.5 used in the paper.
+Notation Description
+$N$ total number of stocks
+$W$ window size of the observation interval
+$D$ total number of features of data
+$T$ total number of trading days
+$K$ codebook size
+$p_{i,t}$ price of stock $i$ at day $t$
+$\mathbf{x}$ observed data, which contains stock historical price data $\mathbf{p}$ and indicators $\mathbf{d}$
+$\mathbf{x}^{\prime}_{ts}$, $\mathbf{x}^{\prime}_{cs}$, reconstructed observed data of the TS and CS modules
+$\hat{\mathbf{y}}_{i,t}$, $\mathbf{y}_{i,t}$ predicted and true return of stock $i$ at day $t$
+$\mathbf{w}_{t}$ the portfolio weight vector for $N$ stocks at trading day $t$
+$a_{i,t}$ action of day $t$ on the stock $i$
+$s_{t}$ state of day $t$
+$\mathbf{z}_{e}^{ts}(\mathbf{x})$, $\mathbf{z}_{e}^{cs}(\mathbf{x})$ latent embeddings of the TS and CS modules
+$\mathbf{z}_{q}^{ts}(\mathbf{x})$, $\mathbf{z}_{q}^{cs}(\mathbf{x})$ quantized embeddings of the TS and CS modules
+$\mathbf{z}_{e}(\mathbf{x})$ factor embeddings
+$\mathbf{z}$ posterior or latent factors
+$k$ total number of selected stocks in PM strategy
+$d$ the max number of stocks to be sold out
+Table A.5. Notations in the paper.
+Appendix B Implement Details
 We implement all experiments by using PyTorch and conduct the experiments on two 80GB memory NVIDIA RTX H100 GPUs.
-The code is provided in the repository  [^4] https://github.com/DVampire/Storm.
+The code is provided in the repository  [^4] https://anonymous.4open.science/r/Storm-2823.
 The input data is organized into batches with dimensions $(B,W,N,D)$, where $B=16$ denotes the batch size.
@@
 The dimension $W=64$ corresponds to the number of historical data days, and $D=152$ represents the total number of features, including the OHLC data, technical indicators, and temporal information.
-For the time-series module, we use an 8-day period for each stock as a patch, meaning the patch size is $(8,1,D)$.
+For the time-series module, we use an 4-day period for each stock as a patch, meaning the patch size is $(4,1,D)$.
 In contrast, for the cross-sectional module, we use $N$ stocks for each day as a patch, so the patch size is $(1,N,D)$.
@@
 Each encoder consists of 4 layers of blocks with 4 attention heads, while each decoder consists of 2 layers of blocks with 8 attention heads.
-The size of both codebooks is selected from the set {256, 512, 1024}, with 512 yielding the best results.
+The size of both codebooks is selected from the set {128, 256, 512, 1024, 2048}, with 512 yielding the best results.
 A consistent embedding dimension of 256 is employed across the encoders, decoders, and codebooks.
@@
 Notably, the default parameters provided by the official implementations are used for baseline models, ensuring a consistent and fair comparison across different approaches.
-5.3. Performance on Factor Quality (for RQ1)
-To demonstrate the effectiveness and quality of the factors, we designed a stock future return prediction task to showcase the factors’ predictive ability and robustness to capture dynamic market changes.
-In the prediction task, STORM shows substantial improvements over six baseline methods across the SP500 and DJ30 datasets.
-Specifically, STORM achieves an average improvement of 16.105% on SP500 and 16.399% on DJ30, indicating the learned latent factors are significantly effective in capturing stock price trends.
-The diversity of factors is crucial in financial investment, as it ensures the model captures the multi-dimensional characteristics of the market, avoiding redundancy or over-reliance on a single factor.
-In the codebook, each embedding represents a cluster center for a category of factor embeddings.
-To analyze the utilization rate of each factor category, we report the frequency of codebook embedding indices in the testing data.
-For clarity, we group every five indices into one interval, as shown in Figure 3(a).
-The results indicate that the distribution of factor usage is relatively even, demonstrating that the model has achieved good diversity in the factor selection process.
-Figure 3. (a) Distribution of counts for codebook embedding indices. (b) the RankIC and RankICIR results across different codebook sizes.
-We experiment with different codebook sizes on the DJ30 dataset to find the optimal balance between factor diversity and representation quality.
-If the codebook is too large, it may lead to sparse factor representation; if too small, it may result in insufficient factor differentiation and information redundancy.
-As shown in Figure 3(b), codebook size of 512 achieves optimal factor prediction performance.
-5.4. Performance on downstream tasks (for RQ2)
-We use three types of metrics: profit, risk-adjusted profit, and risk to evaluate STORM’s investment performance on downstream tasks, i.e., portfolio management and algorithmic trading.
-For profitability, STORM outperforms all the baselines, with average improvements of 106.10% on APY over the best-performing baseline in the PM task and 10.45% over the best baseline methods on three stocks in the AT task.
-For risk resistance, despite the test periods for both datasets, including the COVID-19 event, STORM still achieves the best AVO value, surpassing all comparison methods in the PM task.
-Moreover, considering the return and risk simultaneously, STORM performs the best on ASR with an average improvement of 58.64% over the best baseline result in the PM task, and surpasses the best baseline results across three stocks in the AT task.
-Such an outstanding ability to balance return and risk is significant in real investment practice.
-5.5. Notations
-We provide the main notations used in the paper.
-Notation Description
-$N$ total number of stocks
-$W$ window size of the observation interval
-$D$ total number of features of data
-$T$ total number of trading days
-$K$ codebook size
-$p_{i,t}$ price of stock $i$ at day $t$
-$\mathbf{x}$ observed data, which contains stock historical price data $\mathbf{p}$ and indicators $\mathbf{d}$
-$\mathbf{x}^{\prime}_{ts}$, $\mathbf{x}^{\prime}_{cs}$, reconstructed observed data of the TS and CS modules
-$\hat{\mathbf{y}}_{i,t}$, $\mathbf{y}_{i,t}$ predicted and true return of stock $i$ at day $t$
-$\mathbf{w}_{t}$ the portfolio weight vector for $N$ stocks at trading day $t$
-$a_{i,t}$ action of day $t$ on the stock $i$
-$s_{t}$ state of day $t$
-$\mathbf{z}_{e}^{ts}(\mathbf{x})$, $\mathbf{z}_{e}^{cs}(\mathbf{x})$ latent embeddings of the TS and CS modules
-$\mathbf{z}_{q}^{ts}(\mathbf{x})$, $\mathbf{z}_{q}^{cs}(\mathbf{x})$ quantized embeddings of the TS and CS modules
-$\mathbf{z}_{e}(\mathbf{x})$ factor embeddings
-$\mathbf{z}$ posterior or latent factors
-$k$ total number of selected stocks in PM strategy
-$d$ the max number of stocks to be sold out
-Table 4. Notations in the paper.
-5.6. Ablation Study (for RQ3)
-We conducted ablation experiments with two simplified versions of STORM, namely STORM-w/o-TS and STORM-w/o-CS, which only extract cross-sectional and time-series factors, respectively.
-Specifically, the absence of either time-series or cross-sectional factors significantly reduces the model’s effectiveness across various tasks.
-In the stock return prediction task, STORM outperforms both simplified models, with an average improvement of 21.73% on the RankIC metric and 24.54% on the RankICIR metric, highlighting the enhanced accuracy achieved by incorporating both time-series and cross-sectional factors.
-In the PM and AT tasks, STORM and two simplified models all performed well.
-Notably, STORM-w/o-TS exhibited the lowest AVO among the three models, although its profitability did not surpass that of STORM.
-These findings suggest that methods like FactorVAE (Duan et al., 2022), which focus solely on cross-sectional factors, may struggle to capture the full complexity of the financial market.
-Therefore, integrating both time-series and cross-sectional factors is essential for robust and accurate financial modeling.
-6. Conclusions and Future Work
-In this paper, we propose STORM, a spatio-temporal latent factor model that leverages the dual VQ-VAE framework to learn both time-series and cross-sectional factors.
-STORM ensures the orthogonality and diversity of factors, which are crucial for effective factor selection in financial trading.
-Extensive experiments demonstrate the superiority of STORM over the state-of-the-art methods across a prediction task and two downstream tasks, showcasing its robust performance and effectiveness.
-In the future, we plan to extend STORM by incorporating more side information, e.g., news, to explore the effects of exogenous factors.

### v4

@@
 The extensive experiments demonstrate STORM’s flexibility in adapting to downstream tasks and superior performance over baseline models.
-Code is available in PyTorch [^1] https://anonymous.4open.science/r/Storm-2823.
+Code is available in PyTorch [^1] https://github.com/DVampire/Storm.
 Factor Model, VQVAE, Portfolio Management, Algorithmic Trading
-conference: Proceedings of the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining 2025; August 03–August 07, 2025; Toronto, Canada
+journalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USAbooktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAdoi: 10.1145/3773966.3777972isbn: 979-8-4007-2292-9/2026/02ccs: Information systems Spatial-temporal systemsccs: Applied computing Electronic commerce
 1. Introduction
-In financial trading, factor models are fundamental tools for asset pricing, widely used to predict asset returns.
-These models enhance the accuracy of asset pricing and risk management by identifying a set of key factors that explain excess returns.
-The well-known Fama-French three-factor model (Fama and French, 1992), for example, builds on the Capital Asset Pricing Model (Sharpe, 1964) by introducing additional factors to capture the risk premiums associated with different types of stocks in the market.
-As financial theories have evolved, researchers have continued to discover new factors, allowing for a more comprehensive understanding of market dynamics.
+In financial trading, factor models are fundamental tools for asset pricing and are widely used to predict asset returns.
+These models enhance pricing accuracy and risk management by identifying a set of key factors that explain excess returns.
+The well-known Fama-French three-factor model (fama1992cross) extends the Capital Asset Pricing Model (sharpe1964capital) with additional factors to capture market risk premiums.
+As financial theories evolve, researchers continue to discover new factors, broadening our understanding of market dynamics.
 Figure 1. Overall architecture of STORM
-Recently, we have witnessed the rise of latent factor models (Gu et al., 2021; Duan et al., 2022; Wei et al., 2022, 2023), connecting the factor model with the generative model, the variational autoencoder (VAE).
+Recently, we have witnessed the rise of latent factor models (gu2021autoencoder; duan2022factorvae; wei2022factor; wei2023hirevae), connecting the factor model with the generative model, the variational autoencoder (VAE).
 These VAE-based models describe high-dimensional data (prices) to low-dimensional representations (factors), and learn factors self-adaptively.
@@
   • CH2: Factor Inefficiency.
-In addition to the potential limitations of factors discussed in CH1, factors learned through VAEs exhibit three primary inefficiencies: i) Most latent factor models only focus on cross-sectional factors, neglecting the temporal perspective.
-ii) The continuous, multi-dimensional latent space may allow noise to overshadow meaningful factors.
-iii) The lack of independence among factors may lead to multicollinearity, with the homogeneity weakening the model’s adaptability to varying market conditions and limiting the effectiveness of factors in downstream investment tasks.
+VAE-learned factors suffer from three inefficiencies: i) focusing mainly on cross-sectional factors while neglecting temporal information, ii) allowing noise in continuous latent spaces to overshadow meaningful signals, and iii) lacking independence among factors, which leads to multicollinearity and weak adaptability to varying market conditions.
   • CH3: Lack of Factor Selection.
@@
 In order to address the challenges, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders (VQ-VAE), named STORM, with the architecture shown in Figure 1.
-The model encodes factors as multi-dimensional embeddings, effectively capturing the complexity and nonlinearity inherent in financial data, and enhancing the capacity of factors to account for factor returns (CH1).
+Unlike traditional scalar-valued financial factors with clear economic interpretations, STORM learns high-dimensional latent vectors that, while less interpretable capture complexity and nonlinearity inherent in financial martket (CH1).
 Additionally, we develop a dual VQ-VAE architecture to capture cross-sectional and time-series features, considering both spatial and temporal perspectives [^2] We treat cross-sectional as spatial and time-series as temporal.
@@
 This strategy provides clarity and transparency to the differentiation and selection process of factors (CH3).
-Specifically, our contributions are three-fold:
+Specifically, our contributions are four-fold:
   • We design a dual VQ-VAE architecture that systematically constructs cross-sectional and time-series factors by capturing both spatial and temporal dependencies.
@@
 Factor model has attracted extensive research focus from the finance community.
-Financial experts have carefully selected a wide range of factors from macroeconomic (Sharpe, 1964), market changes (Fama and French, 1992), individual stock fundamentals (Greenblatt, 2010), and other perspectives to price assets in an attempt to capture excess returns from mispricing.
+Financial experts have carefully selected a wide range of factors from macroeconomic (sharpe1964capital), market changes (fama1992cross), individual stock fundamentals (greenblatt2010little), and other perspectives to price assets in an attempt to capture excess returns from mispricing.
 Recently, latent factor models have emerged that improve investment returns by studying latent factors self-adaptively.
-Specifically, (Gu et al., 2021) learns latent factors depend on asset characteristics, (Duan et al., 2022) extracts cross-sectional factors by predicting future returns, and (Wei et al., 2023) learns market-augmented latent factors.
+Specifically, (gu2021autoencoder) learns latent factors depend on asset characteristics, (duan2022factorvae) extracts cross-sectional factors by predicting future returns, and (wei2023hirevae) learns market-augmented latent factors.
 However, these VAE-based latent factor models may suffer from a lack of robustness in representing low signal-to-noise ratio markets, leading to reduced reconstruction and generation capabilities.
@@
 Financial trading has attracted significant attention from both the finance and AI communities.
-Quantitative traders leverage mathematical models and algorithms to automatically identify trading opportunities (Sun et al., 2024).
-Within the FinTech domain, key research areas encompass portfolio management (PM), algorithmic trading (AT), order execution (Fang et al., 2021), market making (Spooner and Savani, 2020), etc.
+Quantitative traders leverage mathematical models and algorithms to automatically identify trading opportunities (sun2024trademaster).
+Within the FinTech domain, key research areas encompass portfolio management (PM), algorithmic trading (AT), order execution (fang2021universal), market making (spooner2020robust), etc.
 In this work, we focus on PM and AT, as they are fundamental to developing intelligent financial decision-making systems.
-PM focuses on optimizing wealth allocation across assets (Li and Hoi, 2014).
-Its development has evolved from simple rule-based methods (Poterba and Summers, 1988) to prediction-based approaches using machine learning (Ke et al., 2017) and deep learning (Qin et al., 2017; Zhao et al., 2022) for predicting returns or price movements.
-Despite recent integration of reinforcement learning (RL) in PM (Zhang et al., 2024b), the latent factor models still predominantly rely on deep learning for market modeling and precise predictions.
+PM focuses on optimizing wealth allocation across assets (li2014online).
+Its development has evolved from simple rule-based methods (poterba1988mean) to prediction-based approaches using machine learning (ke2017lightgbm) and deep learning (qin2017dual; zhao2022stock) for predicting returns or price movements.
+Despite recent integration of reinforcement learning (RL) in PM (zhang2024reinforcement), the latent factor models still predominantly rely on deep learning for market modeling and precise predictions.
 AT, on the other hand, involves executing trades with algorithm-generated signals.
-Like PM, AT has seen a transition from rule-based and supervised learning techniques (Yang et al., 2020) to RL-driven approaches.
-In particular, RL-based methods (Sun et al., 2022) have gained prominence due to their ability to navigate complex sequential decision-making problems, further enhancing trade execution efficiency and adaptability in dynamic market environments.
+Like PM, AT has seen a transition from rule-based and supervised learning techniques (yang2020qlib) to RL-driven approaches.
+In particular, RL-based methods (sun2022deepscalper) have gained prominence due to their ability to navigate complex sequential decision-making problems, further enhancing trade execution efficiency and adaptability in dynamic market environments.
 2.3. Vector Quantized Variational Autoencoder
 Vector Quantized Variational Autoencoder (VQ-VAE) has emerged as a fundamental framework in deep learning research, establishing methodological significance across core domains: representation learning, generative modeling, and time series analysis.
-Initially introduced by Van Den Oord et al., VQ-VAE has showcased its capabilities in processing high-dimensional data across diverse modalities, including images (Razavi et al., 2019), audio (Zhang et al., 2024a), and video (Yan et al., 2021), with a strong emphasis on learning discrete latent representations.
-Particularly noteworthy is the model’s demonstrated proficiency in capturing temporal dependencies, where recent studies like (Talukder et al., 2024) have validated its superior performance in modeling sequential patterns compared to conventional continuous latent space approaches, thereby advancing the development of robust generative architectures specifically optimized for temporal dynamics analysis.
+Initially introduced by van2017neural(van2017neural), VQ-VAE has showcased its capabilities in processing high-dimensional data across diverse modalities, including images (razavi2019generating), audio (zhang2024codebook), and video (yan2021videogpt), with a strong emphasis on learning discrete latent representations.
+Particularly noteworthy is the model’s demonstrated proficiency in capturing temporal dependencies, where recent studies like (talukder2024totem) have validated its superior performance in modeling sequential patterns compared to conventional continuous latent space approaches, thereby advancing the development of robust generative architectures specifically optimized for temporal dynamics analysis.
 3. Preliminaries
 3.1. Factor Model
-The factor model (e.g., Arbitrage Pricing Theory (Ross, 1976)) is defined to study the relationship between expected asset returns and factor exposures, focusing on the cross-sectional differences in expected asset returns.
+The factor model (e.g., Arbitrage Pricing Theory (ross1976arbitrage)) is defined to study the relationship between expected asset returns and factor exposures, focusing on the cross-sectional differences in expected asset returns.
 (1) $\mathbb{E}[R_{i}^{e}]=\alpha_{i}+\beta_{i}\lambda\text{.}$
@@
 Lastly, $\alpha_{i}$ captures pricing error, reflecting any deviations from the model’s predicted returns.
-According to (Duan et al., 2022), the factor model is formulated from a spatial perspective, primarily examining cross-sectional variations in expected returns across different assets, rather than capturing the temporal evolution of individual asset returns.
+According to (duan2022factorvae), the factor model is formulated from a spatial perspective, primarily examining cross-sectional variations in expected returns across different assets, rather than capturing the temporal evolution of individual asset returns.
 Therefore, we expand it from the temporal perspective and introduce the general functional form of the latent factor model:
-(2) $y_{i,t}=\alpha_{i,t}+\sum_{k=1}^{K}\bm{\beta}_{i,t}^{k}\mathbf{z}_{i,t}^{k}+\epsilon_{i,t}\text{,}\vspace{-0.1cm}$
+(2) $y_{i,t}=\alpha_{i,t}+\sum_{k=1}^{K}\bm{\beta}_{i,t}^{k}\mathbf{z}_{i,t}^{k}+\epsilon_{i,t}\text{,}$
 where $y_{i,t}$ denotes the future return of stock $i$ at timestep $t$, $\mathbf{z}_{i,t}$ denotes $K$ factor returns, and $\epsilon_{i,t}$ is the noise.
@@
 To assess the effectiveness and generating ability of factors, we predict and evaluate the stock future returns.
-To show the adaptability of the latent factor model with multiple downstream tasks, we employ two financial trading tasks, i.e., portfolio management (Wang et al., 2021a; Zhang et al., 2024b; Ye et al., 2020) and algorithmic trading (Sun et al., 2022; Liu et al., 2020), to showcase the model’s investment capabilities.
+To show the adaptability of the latent factor model with multiple downstream tasks, we employ two financial trading tasks, i.e., portfolio management (wang2021deeptrader; zhang2024reinforcement; ye2020reinforcement) and algorithmic trading (sun2022deepscalper; liu2020adaptive), to showcase the model’s investment capabilities.
 Observed Data.
@@
   • Portfolio management task aims to construct an optimal portfolio based on the predicted returns to test the profitability of the factor model.
-At each trading day $t$, we obtain the portfolio weight $\mathbf{w}_{t}=[w_{t}^{0},w_{t}^{1},...,w_{t}^{N-1}]\in\mathbb{R}^{N}$ for $N$ stocks.
-Each stock $i$ is assigned a weight $w_{t}^{i}$ representing its portfolio proportion within the portfolio.
-To ensure full capital deployment, the allocation must satisfy the budge constraint $\sum_{i=0}^{N-1}w_{t}^{i}=1$.
+At each trading day $t$, the factor model outputs return predictions $\hat{\mathbf{y}}_{i,t+1}$ for each stock $i$, which are then used by a portfolio allocation strategy, e.g., TopK-Drop, to generate the portfolio weight vector $\mathbf{w}_{t}=[w_{t}^{0},w_{t}^{1},\dots,w_{t}^{N-1}]\in\mathbb{R}^{N}$, where $w_{t}^{i}$ represents the proportion of capital invested in stock $i$.
+To ensure full capital deployment, the allocation must satisfy the budget constraint $\sum_{i=0}^{N-1}w_{t}^{i}=1$.
   • Algorithmic trading task aims to execute buy, hold, and sell actions based on predicted asset states to balance returns and risks.
-We formulate it as a Markov Decision Process (MDP) under reinforcement learning scenario following (Chen et al., 2023; Zhang et al., 2024c).
+We formulate it as a Markov Decision Process (MDP) under reinforcement learning scenario following (chen2023mtrader; zhang2024finagent).
 The MDP is constructed by a 5-tuple $(\mathit{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where $\mathit{S}$ and $\mathcal{A}$ are sets of states and actions respectively, $\mathcal{T}:\mathit{S}\times\mathcal{A}\times\mathit{S}\rightarrow[0,1]$ is the state transition function, $\mathcal{R}:\mathit{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function where $\mathbb{R}$ is a continuous set of possible rewards, and $\gamma\in[0,1)$ is the discount factor.
+The latent factor embeddings $\mathbf{Z}$ generated by the factor model are integrated into the observation set that constitutes the state space $\mathcal{S}$, enabling the agent to incorporate predictive market signals.
 The goal is to learn a policy $\mathcal{\pi}:\mathit{S}\rightarrow\mathcal{A}$ that maximizes the expected discounted cumulative reward $\mathbb{E}[\sum^{\infty}_{t=0}\gamma^{t-1}r_{t}]$.
@@
 3.3. Vector Quantized Variational Autoencoder
-VQ-VAE(Van Den Oord et al., 2017) is different from VAE (Kingma and Welling, 2013) as it quantizes the observed data into a discrete token sequence.
+VQ-VAE(van2017neural) is different from VAE (kingma2013auto) as it quantizes the observed data into a discrete token sequence.
 The VQ-VAE contains three parts, encoder, decoder, and discrete codebook, denoted by $\phi_{\text{enc}}$ with parameter $\theta_{\text{enc}}$, $\phi_{\text{dec}}$ with parameter $\theta_{\text{dec}}$ and $\mathcal{C}=\{(k,e_{k}\in R^{H})\}_{k=1}^{K}$ with parameter $\{e_{k}\in R^{H}\}_{k=1}^{K}$ respectively.
@@
 Given an input $x$, the encoder $\phi_{\text{enc}}$ firstly encodes it into a set of continuous feature vectors $Z=\phi_{\text{enc}}(x)$ and quantization operation $q(\cdot)$ passes each feature vector $z\in Z$ through the discretization bottleneck following by selecting the nearest neighbor embedding in the codebook $\mathcal{C}$ as its discrete code sequence $Q_{z}$.
-(3) $Q_{z}=q(z,\mathcal{C})=c,\text{where }c=\arg\min_{k\in[0,K-1]}||z-e_{k}||_{2}^{2}\text{.}\vspace{-0.1cm}$
+[…]
 Then, the quantized feature for $z$ denoted as $Z^{q}_{z}$, is obtained through $Z^{q}_{z}=e_{Q_{z}}$.
@@
 Time-series features capture the dynamics of a stock over time, while cross-sectional features reveal the correlations among different stocks.
-To achieve this, we partition data into patches (Nie et al., 2022) within the TS and CS modules.
-In the TS module, the observed data $\mathbf{x}$ is divided along the stock number dimension, with each patch containing data for one stock over $p$ days.
+To achieve this, we partition data into patches (nie2022time) within the TS and CS modules.
+In the TS module, the observed data $\mathbf{x}$ is divided along the stock number dimension, with each patch containing data for one stock over $p$ days (as illustrated in Figure 2).
 Similarly, in the CS module, the observed data $\mathbf{x}$ is divided along the time axis, with each patch containing features of all stocks for a single trading day.
@@
 4.1.2. TS and CS Encoders
-Considering the Transformer’s (Vaswani et al., 2017) remarkable ability to capture both local and global dependencies via self-attention and patching strategies, it is particularly well-suited for long-term time-series forecasting (Nie et al., 2022).
+Considering the Transformer’s (vaswani2017attention) remarkable ability to capture both local and global dependencies via self-attention and patching strategies, it is particularly well-suited for long-term time-series forecasting (nie2022time).
 Therefore, we employ stacked transformer blocks as the encoders in both the TS and CS modules to capture complex time-series and cross-sectional patterns.
 This architecture enhances the accuracy and robustness of feature extraction, ultimately improving forecasting performance.
-While transformers are widely applicable across domains such as natural language processing (NLP) and computer vision (CV), STORM’s flexible encoder-decoder framework allows for the integration of alternative architectures, such as long short-term memory (LSTM) (Wang et al., 2021b) and temporal convolutional networks (TCN) (Yu and Koltun, 2015).
+While transformers are widely applicable across domains such as natural language processing (NLP) and computer vision (CV), STORM’s flexible encoder-decoder framework allows for the integration of alternative architectures, such as long short-term memory (LSTM) (wang2021clvsa) and temporal convolutional networks (TCN) (yu2015multi).
+An encoder–decoder configuration comprising 4 and 2 blocks, respectively, each with 4 and 8 attention heads, remains relatively lightweight, avoiding excessive computational cost and a substantial increase in model complexity.
 4.1.3. Codebook Construction and Optimization
-The construction of the codebook involves a learnable embedding space designed to quantize the encoded features into discrete representations (Van Den Oord et al., 2017).
+The construction of the codebook involves a learnable embedding space designed to quantize the encoded features into discrete representations (van2017neural).
 During training, the continuous latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$ are quantized into discrete space by mapping each vector to its nearest codebook entry, $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$, based on minimizing the Euclidean distance.
@@
 iii) Noise Reduction: By filtering out irrelevant or redundant information during factor selection, VQ-VAE mitigates noise, contributing to improved model robustness and more reliable financial insights.
-To encourage equal usage of the codebook vectors (Baevski et al., 2020), we incorporate a diversity loss to ensure balanced representations across the codebook.
+To encourage equal usage of the codebook vectors (baevski2020wav2vec), we incorporate a diversity loss to ensure balanced representations across the codebook.
 It is achieved by maximizing the entropy of the averaged softmax distribution over vectors for each codebook $\bar{p}_{g}$:
-(4) $\mathcal{L}_{\text{div}}=\frac{1}{GK}\sum_{g=1}^{G}\sum_{k=1}^{K}\bar{p}_{g,k}\log{\bar{p}_{g,k}}\text{,}\vspace{-0.1cm}$
+(4) $\mathcal{L}_{\text{div}}=\frac{1}{GK}\sum_{g=1}^{G}\sum_{k=1}^{K}\bar{p}_{g,k}\log{\bar{p}_{g,k}}\text{,}$
 where $G$ represents the number of codebooks, set 2 in STORM. The diversity constraint helps enhance representational capacity, leading to better coverage of the input space.
 However, the lack of independence among factors can introduce multicollinearity, making the model prone to overfitting and reducing its robustness across different market conditions.
-To address this issue, we impose an orthogonality constraint on the codebook, as recent studies (Shin et al., 2023) have shown that enforcing orthogonality allows discretized codes to maintain translation equivariance.
+To address this issue, we impose an orthogonality constraint on the codebook, as recent studies (shin2023exploration) have shown that enforcing orthogonality allows discretized codes to maintain translation equivariance.
 The orthogonality loss is:
-(5) $\mathcal{L}_{\text{ortho}}=\frac{1}{K^{2}}\left|\left|\ell_{2}(\mathbf{e})^{\top}\ell_{2}(\mathbf{e})-I_{K}\right|\right|_{F}^{2}\text{,}\vspace{-0.1cm}$
+[…]
 where $I_{k}\in\mathbb{R}^{K\times K}$ is the identity matrix, $\mathbf{e}$ is a surrogate for two codebook embeddings $\mathbf{ts}$ and $\mathbf{cs}$, $\ell_{2}(\mathbf{e})$ denotes $L_{2}$-normalized embeddings, and $\left|\left|\cdot\right|\right|_{F}$ is the Frobenius norm.
@@
 Finally, the encoders, decoders, and codebooks are jointly optimized by minimizing the following loss objectives:
-(6) $\mathcal{L}_{1}=$ $\lambda_{\text{ortho}}\mathcal{L}_{\text{ortho}}+\lambda_{\text{div}}\mathcal{L}_{\text{div}}+\left|\left|\mathbf{x}-\mathbf{x}^{\prime}_{ts}\right|\right|^{2}_{2}+\left|\left|\mathbf{x}-\mathbf{x}^{\prime}_{cs}\right|\right|^{2}_{2}$
-$+$ $\left|\left|sg[\mathbf{z}_{e}^{ts}(\mathbf{x})]-\mathbf{z}_{q}^{ts}(\mathbf{x})\right|\right|^{2}_{2}+\left|\left|sg[\mathbf{z}_{q}^{ts}(\mathbf{x})]-\mathbf{z}_{e}^{ts}(\mathbf{x})\right|\right|^{2}_{2}$
-$+$ $\left|\left|sg[\mathbf{z}_{e}^{cs}(\mathbf{x})]-\mathbf{z}_{q}^{cs}(\mathbf{x})\right|\right|^{2}_{2}+\left|\left|sg[\mathbf{z}_{q}^{cs}(\mathbf{x})]-\mathbf{z}_{e}^{cs}(\mathbf{x})\right|\right|^{2}_{2}\text{,}$
+[…]
+[…]
+[…]
 where $sg[\cdot]$ is a stop-gradient operator.
@@
 Cross-Attention Mechanism.
-In order to fuse the features in the TS and CS modules at a fine-grained level, we leverage a multi-scale encoder to combine information between two patches and produce stronger fusion features for factor generations, motivated by CrossViT (Chen et al., 2021).
-The architecture incorporates cross-attention layers, balancing both accuracy and computational efficiency, ensuring that factors derived from the combined features are more representative of market dynamics.
+In order to fuse the features in the TS and CS modules at a fine-grained level, we leverage a multi-scale encoder to combine information between two temporal and spatial patterns and produce stronger fusion features for factor generations, motivated by CrossViT (chen2021crossvit).
+The architecture incorporates cross-attention layers that explicitly learn correlations between two patterns, enabling the encoder to emphasize mutually informative regions while filtering out noise.
+This design balances accuracy and computational efficiency, ensuring the fused features yield richer and more discriminative signals for factor generation and that the resulting factors are more representative of market dynamics.
 Contrastive Learning.
-To further enhance semantic consistency between TS and CS features, we integrate contrastive learning, drawing inspiration from its success in aligning multimodal representations (Li et al., 2023; Radford et al., 2021).
-Within the factor module, a contrastive loss encourages similar features to cluster closely while pushing apart dissimilar ones in the embedding space.
+To further enhance semantic consistency between TS and CS features, we integrate contrastive learning, drawing inspiration from its success in aligning multimodal representations (li2023blip; radford2021learning).
+A contrastive loss is introduced within the factor module to encourage features from TS and CS modules that represent similar market conditions to cluster closely in the embedding space, while pushing apart features corresponding to different conditions.
+This dual-stream alignment improves the coherence of the factor space and enhances the model’s ability to generalize across diverse market scenarios.
 4.3.2. Prior-Posterior Learning
 Through multi-scale encoder and contrastive learning layer, the CS and TS latent features are fused and aligned.
-In order to retain the codebook’s categorical properties for factors, we incorporate the codebook embeddings $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$ as extra [CLS]-tokens (Dosovitskiy et al., 2020), which are appended to the latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$, respectively.
+In order to retain the codebook’s categorical properties for factors, we incorporate the codebook embeddings $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$ as extra [CLS]-tokens (dosovitskiy2020image), which are appended to the latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$, respectively.
 Then, two latent features concat together to get the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$.
-Given the inherent volatility and complexity of the stock market, bridging the gap between noisy market data and an effective factor model remains a significant challenge (Duan et al., 2022).
+Given the inherent volatility and complexity of the stock market, bridging the gap between noisy market data and an effective factor model remains a significant challenge (duan2022factorvae).
 Therefore, we introduce the prior-posterior structure to predict future returns and optimize latent factor embeddings.
 In the training stage, the posterior layer estimates posterior distribution of factor expected returns $\mathbf{z}_{post}$ from the true future stock returns $\mathbf{y}$ and the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$:
-(7) $[\mu_{post},\sigma_{post}]=\phi_{FE}(\mathbf{y},\mathbf{z}_{e}(\mathbf{x}))\text{,}$
+(7) $\mu_{post},\sigma_{post}]=\phi_{FE}(\mathbf{y},\mathbf{z}_{e}(\mathbf{x}))\text{,}$
 and $\mathbf{z}_{post}$ follows the independent Gaussian distribution.
 In the inference stage, the prior layer only use latent factor embeddings without any future information leakage to calculate prior distribution of factor expected returns $\mathbf{z}_{prior}$:
-(8) $[\mu_{prior},\sigma_{prior}]=\phi_{FP}(\mathbf{z}_{e}(\mathbf{x}))\text{.}$
+(8) $\mu_{prior},\sigma_{prior}]=\phi_{FP}(\mathbf{z}_{e}(\mathbf{x}))\text{.}$
 The return predictor computes future returns based on the factor expected return $\mathbf{z}$:
@@
 4.4. Downstream Tasks
-Existing research on latent factor models (Duan et al., 2022; Wei et al., 2023) primarily focuses on predicting future returns by analyzing cross-sectional characteristics of market stocks, which means the profitability of the factor model in quantitative trading is typically evaluated through portfolio management tasks.
-In contrast, STORM captures both the cross-sectional features among stocks and the time-series features of individual stocks, enabling it to excel in both PM and single-asset AT tasks.
+Existing research on latent factor models (duan2022factorvae; wei2023hirevae) primarily focuses on predicting future returns by analyzing cross-sectional characteristics of market stocks, which means the profitability of the factor model in financial investment is typically evaluated through PM tasks.
+In contrast, STORM captures the cross-sectional features among stocks and the time-series features of individual stocks, enabling it to excel in both PM and single-asset AT tasks.
 4.4.1. Portfolio Management
-To evaluate the performance of the STORM and compare it with baseline methods, we follow the paradigm proposed by FactorVAE (Duan et al., 2022), focusing on return prediction and the PM task.
-Specifically, we utilize the factor decoder network to generate stock future returns $\mathbf{\hat{y}}$, and then apply the TopK-Drop strategy [^3] https://qlib.readthedocs.io/en/latest/component/strategy.html to backtest the factor model.
-The TopK-Drop strategy constructs a daily portfolio consisting of top $k$ stocks based on the return predictions.
-Considering the transaction costs in the real stock market, the strategy sets the turnover constraint, which ensures that more than $k-d$ stocks overlap between the portfolios $\mathbf{w}_{t}$ and $\mathbf{w}_{t+1}$ on consecutive trading days.
+To evaluate the performance of the STORM and compare it with baseline methods, we follow the paradigm proposed by FactorVAE (duan2022factorvae), focusing on return prediction and the PM task.
+Specifically, we utilize the factor decoder network to generate stock future returns $\mathbf{\hat{y}}$, and then apply the TopK-Drop strategy [^3] https://qlib.readthedocs.io/en/latest/component/strategy.html, which constructs a daily portfolio by selecting the top $k$ stocks based on predicted returns, to backtest the factor model.
+Considering transaction costs in real markets, there is a turnover constraint that limits the number of changes in the portfolio composition between consecutive trading days.
+Specifically, when moving from $\mathbf{w}_{t}$ to $\mathbf{w}_{t+1}$, at most $d$ stocks can be replaced, where a replacement is defined as a stock whose weight changes from 0 (not held) to a positive value (newly bought) or from a positive value to 0 (fully sold).
+This constraint ensures that at least kdk-d stocks remain in the portfolio, reducing excessive turnover.
 4.4.2. Algorithmic trading
@@
 In the AT task, since it involves trading a single stock, the features of other stocks are excluded.
-We use the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) to optimize the policy.
+We use the Proximal Policy Optimization (PPO) algorithm (schulman2017proximal) to optimize the policy.
+In each iteration, the agent interacts with the environment to collect trajectories, computes advantage estimates, and updates the policy parameters via PPO’s clipped surrogate objective.
+This optimization process ensures the stability and efficiency of policy updates, helping the agent gradually improve the decision-making ability and achieve better trading strategies.
+Table 1. Comparison of baselines across different tasks
+Task Market ML & DL-based RL-based Factor model Ours
+B&H LGBM (yang2020qlib) LSTM (yu2018forecasting) Transformer (yang2020qlib) SAC (haarnoja2018soft) PPO (schulman2017proximal) DQN (mnih2013playing) CAFactor (gu2021autoencoder) FactorVAE (duan2022factorvae) HireVAE (wei2023hirevae) STORM
+Prediction
+Portfolio Management
+Algorithmic Trading
+  • * indicates that the model is applicable, while signifies that it is NOT applicable to the corresponding task.
 5. Experiment
 In this section, we evaluate the proposed STORM on real stock markets and conduct extensive experiments to address the following research questions.
-RQ1: How to evaluate the effectiveness of STORM’s learned factors?
-RQ2: How does STORM perform on downstream tasks?
-RQ3: How do the key components contribute to the performance of STORM?
+RQ1:
+How does STORM perform on downstream tasks?
+RQ2:
+How to evaluate the effectiveness of STORM’s learned factors?
+RQ3:
+How do the key components contribute to the performance of STORM?
 5.1. Experiment Settings
-Dataset.
+5.1.1. Datasets
 We conduct experiments on two U.S.
-stock markets, SP500 and DJ30, using stock daily data that includes technical features based on Alpha158 (Yang et al., 2020).
+stock markets, SP500 and DJ30, using stock daily data that includes technical features based on Alpha158 (yang2020qlib).
 Both datasets span 16 years, from 2008-04-01 to 2024-03-31, encompassing global conditions, e.g., the 2007-2008 financial crisis and COVID-19.
 Datasets are chronologically divided into non-overlapping training (from 2008-04-01 to 2021-03-31) and test (from 2021-04-01 to 2024-03-31) sets.
-Metrics.
+5.1.2. Metrics
 We compare STORM and baselines in terms of 6 financial metrics across the PM and AT tasks.
@@
   • RankIC is a ranking metric in finance, which measures the correlation between the predicted rankings and the actual returns. It is defined as:
-    (10) $\text{RankIC}_{s}=\frac{1}{N}\frac{(r_{\hat{y}_{s}}-\text{mean}(r_{\hat{y}_{s}}))^{T}(r_{y_{t}}-\text{mean}(r_{y_{s}}))}{\text{std}(r_{\hat{y}_{s}})\cdot\text{std}(r_{y_{s}})}$
+    $\text{RankIC}_{s}=\frac{1}{N}\frac{(r_{\hat{y}_{s}}-\text{mean}(r_{\hat{y}_{s}}))^{T}(r_{y_{t}}-\text{mean}(r_{y_{s}}))}{\text{std}(r_{\hat{y}_{s}})\cdot\text{std}(r_{y_{s}})}$
     $\text{RankIC}=\frac{1}{T_{\text{test}}}\sum_{s=1}^{T_{\text{test}}}\text{RankIC}_{s}\text{,}$
     where $T_{test}$ denotes the number of trading days in the test range, $r_{y_{s}}$ and $r_{\hat{y}_{s}}$ represent the true and predicted ranks of stocks on the trading day $s$, respectively.
-  • RankICIR is the information ratio of RankIC, which measures the stability of prediction, $\text{RankICIR}=\frac{\text{mean}(\text{RankIC}_{s})}{\text{std}(\text{RankIC}_{s})}$.
-Baselines.
-As shown in Table 2, we compare STORM with nine methods across one prediction task and two downstream tasks.
-The baseline methods can be divided into three categories: ML & DL-based models, Factor models, and RL-based models.
-We evaluate STORM against ML&DL models and factor models in prediction and PM tasks, and against ML&DL models and RL methods in the AT task.
-Moreover, STORM is a versatile all-rounder that supports three tasks, unlike other methods specialized in specific tasks.
-The following will provide a brief introduction to each method:
-  • Market
-    • Buy-and-Hold (B&H) involves holding assets for an extended period, regardless of short-term market fluctuations, assuming that long-term returns will be more favorable.
-  • ML&DL-based
-    • LGBM (Yang et al., 2020) uses a series of tree models to predict price fluctuations and provide buy and sell signals.
-    • LSTM (Yu and Li, 2018) utilizes long short-term memory to improve the accuracy of price predictions.
-    • Transformer (Yang et al., 2020) models leverage self-attention mechanisms to enhance the precision of price forecasts.
-  • RL-based
-    • SAC (Haarnoja et al., 2018) is an off-policy actor-critic algorithm that optimizes trading strategies using entropy regularization and soft value functions in continuous action spaces.
-    • PPO (Schulman et al., 2017) updates trading policies iteratively to balance exploration and exploitation, ensuring stability and sample efficiency.
-    • DQN (Mnih et al., 2013) uses deep neural networks to approximate the action-value function and make trading decisions from market data.
-  • Factor-model
-    • CAFactor (Gu et al., 2021) is a state-of-the-art model that learns latent factors depending on asset characteristics.
-    • FactorVAE (Duan et al., 2022) is a state-of-the-art VAE-based latent factor model that learns optimal latent stock factors.
-    • HireVAE (Wei et al., 2023) is a state-of-the-art VAE-based latent factor model that identifies factors between market situation and stock-wise.
-Models SP500 DJ30
-[…]
-Future Return Prediction in Portfolio Management Task
-LightGBM 0.027 0.274 0.031 0.272
-± 0.006 ± 0.084 ± 0.005 ± 0.049
-LSTM 0.034 0.333 0.031 0.329
-± 0.006 ± 0.042 ± 0.004 ± 0.056
-Transformer 0.035 0.340 0.033 0.343
-± 0.007 ± 0.078 ± 0.005 ± 0.045
-CAFactor 0.037 0.356 0.040 0.380
-± 0.005 ± 0.084 ± 0.003 ± 0.043
-FactorVAE 0.052 0.543 0.056 0.520
-± 0.010 ± 0.122 ± 0.012 ± 0.081
-HireVAE 0.057 0.558 0.058 0.563
-± 0.006 ± 0.058 ± 0.006 ± 0.053
-STORM 0.063 0.679 0.066 0.670
-± 0.020 ± 0.169 ± 0.038 ± 0.303
-STORM-w/o-TS 0.051 0.514 0.055 0.560
-± 0.016 ± 0.141 ± 0.016 ± 0.105
-STORM-w/o-CS 0.053 0.514 0.053 0.586
-± 0.013 ± 0.126 ± 0.017 ± 0.143
-Improvement(%)11footnotemark: 1 10.526 21.685 13.793 19.005
-  • $1$
-    Improvement of STORM over the best-performing baselines.
-Table 1. Results of all models on two metrics, i.e., RankIC and RankICIR (mean ± range, computed across 10 runs).
-Task Market ML & DL-based RL-based Factor model Ours
-B&H LGBM (Yang et al., 2020) LSTM (Yu and Li, 2018) Transformer (Yang et al., 2020) SAC (Haarnoja et al., 2018) PPO (Schulman et al., 2017) DQN (Mnih et al., 2013) CAFactor (Gu et al., 2021) FactorVAE (Duan et al., 2022) HireVAE (Wei et al., 2023) STORM
-Prediction
-Portfolio Management
-Algorithmic Trading
-  • * indicates that the model is applicable, while signifies that it is NOT applicable to the corresponding task.
-Table 2. Comparison of baselines across different tasks
-Implementation Details.
-All experiments are implemented using PyTorch and conducted on two NVIDIA RTX H100 GPUs with 80GB of memory each.
-The implementations of the ML & DL methods are based on Qlib (Yang et al., 2020).
-The input data is organized into batches with dimensions corresponding to historical days, stocks, and features, ensuring compatibility across tasks.
-As for other baselines, we adopt the default settings in their public implementations.
-Each experiment is repeated multiple times with different random seeds, and the mean and variance of the metrics are reported.
-Detailed information regarding the network architecture of STORM, as well as the optimizers and hyperparameters used for the two downstream tasks, is provided in Appendix B.
+  • RankICIR is the information ratio of RankIC, which measures the stability of prediction,
+    $\text{RankICIR}=\frac{\text{mean}(\text{RankIC}_{s})}{\text{std}(\text{RankIC}_{s})}\text{.}$
+Table 2. Portfolio management task results of all models across six metrics (mean ± range, computed across 10 runs).
 SP500 Dataset
@@
 ± 0.020 ± 0.068 ± 0.065 ± 0.114 ± 0.042 ± 0.013
-Transformer 0.076 1.244 0.389 0.433 0.198 0.174
+Transformer 0.076 1.244 0.389 0.433 0.19811footnotemark: 1 0.174
 ± 0.028 ± 0.098 ± 0.165 ± 0.161 ± 0.057 ± 0.004
@@
 ± 0.029 ± 0.104 ± 0.180 ± 0.189 ± 0.048 ± 0.007
-STORM 0.177 1.641 1.193 0.817 0.692 0.072
-± 0.103 ± 0.457 ± 0.280 ± 0.071 ± 0.091 ± 0.011
-STORM-w/o-TS 0.084 1.278 1.055 0.702 0.632 0.068
-± 0.060 ± 0.204 ± 0.271 ± 0.062 ± 0.107 ± 0.003
-STORM-w/o-CS 0.080 1.264 0.971 0.698 0.684 0.068
-± 0.052 ± 0.176 ± 0.192 ± 0.065 ± 0.071 ± 0.005
-Improvement(%) 124.05 30.65 195.30 77.61 - 60.47
+STORM 0.18822footnotemark: 2 1.683 1.189 1.052 0.166 0.171
+± 0.055 ± 0.226 ± 0.661 ± 0.329 ± 0.050 ± 0.020
+STORM-w/o-TS 0.090 1.300 0.503 0.570 0.181 0.175
+± 0.074 ± 0.277 ± 0.390 ± 0.339 ± 0.043 ± 0.025
+STORM-w/o-CS 0.089 1.294 0.623 0.592 0.146 0.167
+± 0.045 ± 0.163 ± 0.385 ± 0.252 ± 0.073 ± 0.020
+Improvement(%)33footnotemark: 3 137.97 34.00 194.31 128.70 26.26 2.91
 DJ30 Dataset
@@
 ± 0.013 ± 0.043 ± 0.067 ± 0.085 ± 0.033 ± 0.002
-CAFactor 0.059 1.186 0.233 0.38 0.252 0.153
+CAFactor 0.059 1.186 0.233 0.382 0.252 0.153
 ± 0.015 ± 0.050 ± 0.058 ± 0.092 ± 0.021 ± 0.001
@@
 ± 0.037 ± 0.132 ± 0.253 ± 0.256 ± 0.057 ± 0.012
-STORM 0.143 1.501 0.839 0.680 0.644 0.057
-± 0.078 ± 0.321 ± 0.213 ± 0.080 ± 0.102 ± 0.004
-STORM-w/o-TS 0.071 1.244 0.757 0.596 0.634 0.056
-± 0.134 ± 0.509 ± 0.375 ± 0.144 ± 0.159 ± 0.004
-STORM-w/o-CS 0.077 1.266 0.764 0.607 0.632 0.057
-± 0.156 ± 0.612 ± 0.203 ± 0.18 ± 0.118 ± 0.004
-Improvement(%) 88.16 20.47 138.35 41.67 - 63.40
-Table 3. Portfolio management task results of all models across six metrics (mean ± range, computed across 10 runs).
+STORM 0.148 1.517 1.396 1.052 0.108 0.140
+± 0.046 ± 0.188 ± 0.679 ± 0.297 ± 0.026 ± 0.014
+STORM-w/o-TS 0.079 1.259 0.621 0.603 0.127 0.142
+± 0.057 ± 0.192 ± 0.425 ± 0.379 ± 0.038 ± 0.008
+STORM-w/o-CS 0.073 1.236 0.533 0.566 0.138 0.140
+± 0.051 ± 0.169 ± 0.400 ± 0.343 ± 0.034 ± 0.017
+Improvement(%) 94.74 21.75 296.59 119.17 52.00 8.50
+  • 11
+    Underline indicates the best-performing baseline method result;
+  • 22
+    Bold indicates the best performance among our proposed models (including variants);
+  • 33
+    Improvement of STORM over the best-performing baselines.
+Table 3. Algorithmic trading task results on all models across six standard metrics (averaged over 5 runs).
 Models AAPL JPM IBM INTC MSFT
@@
 STORM 1.346 0.282 0.016 1.245 0.299 0.014 2.107 0.137 0.009 1.189 0.227 0.015 1.672 0.210 0.015
-± 0.162 ± $0$ 0.001 ± 0.106 ± 0.016 ± 0.001 ± 0.302 ± 0.033 ± 0.001 ± 0.386 ± 0.055 ± 0.002 ± 0.201 ± 0.067 ± 0.002
+± 0.162 ± 0.000 ± 0.001 ± 0.106 ± 0.016 ± 0.001 ± 0.302 ± 0.033 ± 0.001 ± 0.386 ± 0.055 ± 0.002 ± 0.201 ± 0.067 ± 0.002
 STORM-w/o-TS 1.199 0.262 0.017 1.587 0.194 0.011 1.868 0.131 0.008 0.580 0.344 0.017 1.553 0.244 0.015
-± 0.300 ± 0.052 ± 0.001 ± 0.296 ± 0.038 ± 0.001 ± 0.956 ± 0.077 ± 0.003 ± 0.229 ± 0.061 ± $0$ 0.187 ± 0.093 ± 0.004
+± 0.300 ± 0.052 ± 0.001 ± 0.296 ± 0.038 ± 0.001 ± 0.956 ± 0.077 ± 0.003 ± 0.229 ± 0.061 ± 0.000 ± 0.187 ± 0.093 ± 0.004
 STORM-w/o-CS 1.027 0.229 0.015 0.858 0.327 0.013 1.397 0.177 0.010 0.914 0.312 0.018 1.401 0.295 0.016
@@
 Improvement(%) 27.704 4.184 - 19.144 8.491 - 15.642 16.561 11.111 95.238 - - 2.076 - -
-Table 4. Algorithmic trading task results on all models across six standard metrics (averaged over 5 runs).
-5.2. Factor Quality (for RQ1)
-To demonstrate the effectiveness and quality of the factors, we designed a stock future return prediction task to showcase the factors’ predictive ability and robustness to capture dynamic market changes.
-In the prediction task, STORM shows substantial improvements over six baseline methods across the SP500 and DJ30 datasets.
-Specifically, in terms of RankIC and RankICIR, STORM achieves an average improvement of 16.105% on SP500 and 16.399% on DJ30, as shown in Table  1.
-These results highlight the learned latent factors are significantly effective in capturing stock price trends.
+5.1.3. Baselines
+As shown in Table 1, we compare STORM with nine methods across one prediction task and two downstream tasks.
+The baseline methods fall into three categories: ML & DL-based models, Factor models, and RL-based models.
+We evaluate STORM against ML&DL models and factor models in prediction and PM tasks, and against ML&DL models and RL methods in the AT task.
+Moreover, STORM is a versatile all-rounder that supports three tasks, unlike other methods specialized in specific tasks.
+  • Market
+    • Buy-and-Hold (B&H) involves holding assets for an extended period, regardless of short-term market fluctuations, assuming that long-term returns will be more favorable.
+  • ML&DL-based
+    • LGBM (yang2020qlib) uses a series of tree models to predict price fluctuations and provide buy and sell signals.
+    • LSTM (yu2018forecasting) utilizes long short-term memory to improve the accuracy of price predictions.
+    • Transformer (yang2020qlib) models leverage self-attention mechanisms to enhance the precision of price forecasts.
+  • RL-based
+    • SAC (haarnoja2018soft) is an off-policy actor-critic algorithm using entropy regularization and soft value functions.
+    • PPO (schulman2017proximal) updates trading policies iteratively to balance exploration and exploitation, ensuring stability and efficiency.
+    • DQN (mnih2013playing) approximates the action-value function and make trading decisions from market data with a deep Q-network.
+  • Factor-model
+    • CAFactor (gu2021autoencoder) learns latent factors from asset features.
+    • FactorVAE (duan2022factorvae) is a state-of-the-art VAE-based latent factor model that learns optimal latent stock factors.
+    • HireVAE (wei2023hirevae) is a state-of-the-art VAE-based latent factor model capturing market and inter-stock factors.
+5.1.4. Implementation Details
+All experiments are implemented using PyTorch and conducted on two NVIDIA RTX H100 GPUs with 80GB of memory each.
+The input data is organized into batches with dimensions $(B,W,N,D)$, where $B=16$ is the batch size, $N$ is the number of stocks (28 for DJ30, 408 for SP500), $W=64$ denotes historical days, and $D=152$ is the total number of features, including the OHLC data and technical indicators.
+The TS module uses patches of size $(4,1,D)$ (4-day sequences per stock), whereas the CS module uses $(1,N,D)$ (all stocks on a given day).
+Encoders (4 layers, 4 heads) and decoders (2 layers, 8 heads) are standard Transformer blocks with a shared embedding size of 256.
+Codebook sizes are searched in 128,256,512,1024,2048{128,256,512,1024,2048}, with 512 performing best.
+[…]
+For downstream tasks, we use Qlib’s [^4] https://github.com/microsoft/qlib TopK-Drop portfolio strategy (k=5k{=}5, d=3d{=}3) and PPO for trading.
+The pretrained TS/CS factors serve as states; all CS factors are retained while only TS factors relevant to traded stocks are used.
+The action space is buy, hold, sell.
+Rewards follow $r=\frac{v_{\text{post}}-v_{\text{pre}}}{v_{\text{pre}}}$ with $v=p\times m+c$, starting from $10^{6}$ initial cash and transaction cost $10^{-4}$.
+The implementations of the ML & DL methods are based on Qlib (yang2020qlib).
+As for other baselines, we adopt the default settings in their public implementations.
+Each experiment is repeated multiple times with different random seeds, and the mean and variance of the metrics are reported.
+5.2. Performance on Investment Tasks (for RQ1)
+We use three types of metrics: profit, risk, and risk-adjusted profit to evaluate STORM’s investment performance on downstream tasks, i.e., portfolio management and algorithmic trading.
+The evaluation results are presented in Table 2 and Table 3.
+For profitability, STORM outperforms all the baselines, with average improvements of 116.36% on APY over the best-performing baseline in the PM task and 43.64% over the best baseline methods on five stocks in the AT task.
+These results underscore the model’s capability to extract meaningful factors, and optimize decision-making for enhanced financial performance.
+For risk resistance, although STORM does not outperform the best baseline methods in terms of MDD and AVO on the INTC and MSFT datasets for the trading task, this can be attributed to the extreme price fluctuations during the COVID-19 event.
+Both stocks experienced a sharp decline of over 30%, followed by a subsequent rebound.
+Given STORM’s strong profitability capability, its trading strategy tends to aggressively capture trends, which naturally leads to larger drawdowns and increased volatility as a trade-off for higher potential returns.
+Moreover, considering the return and risk simultaneously, STORM performs the best on ASR and CR with an average improvement of 123.94% and 245.45% respectively, over the best baseline result in the PM task, and surpasses the best baseline results across three stocks in the AT task.
+Such an outstanding ability to balance return and risk is particularly valuable in real-world investment, where achieving high returns while effectively managing risk is crucial for long-term financial success.
+Table 4. Factor quality evaluation task results on RankIC and RankICIR (mean ± range, computed across 10 runs).
+Models SP500 DJ30
+[…]
+Future Return Prediction in Portfolio Management Task
+LightGBM 0.027 0.274 0.031 0.272
+± 0.006 ± 0.084 ± 0.005 ± 0.049
+LSTM 0.034 0.333 0.031 0.329
+± 0.006 ± 0.042 ± 0.004 ± 0.056
+Transformer 0.035 0.340 0.033 0.343
+± 0.007 ± 0.078 ± 0.005 ± 0.045
+CAFactor 0.037 0.356 0.040 0.380
+± 0.005 ± 0.084 ± 0.003 ± 0.043
+FactorVAE 0.052 0.543 0.056 0.520
+± 0.010 ± 0.122 ± 0.012 ± 0.081
+HireVAE 0.057 0.558 0.058 0.563
+± 0.006 ± 0.058 ± 0.006 ± 0.053
+STORM 0.062 0.673 0.065 0.668
+± 0.018 ± 0.155 ± 0.038 ± 0.287
+STORM-w/o-TS 0.053 0.513 0.055 0.563
+± 0.017 ± 0.145 ± 0.015 ± 0.127
+STORM-w/o-CS 0.054 0.522 0.053 0.559
+± 0.014 ± 0.118 ± 0.016 ± 0.156
+Improvement(%)11footnotemark: 1 8.772 20.609 12.069 18.650
+  • 11
+    Improvement of STORM over the best-performing baselines.
+5.3. Factor Quality (for RQ2)
+To demonstrate the effectiveness and quality of the factors, we designed a stock future return prediction task.
+The goal is not to pursue time-series forecasting accuracy, but to evaluate the predictive quality and robustness of the factors in capturing dynamic market patterns.
+In this task, STORM shows substantial improvements over six methods across the SP500 and DJ30 datasets.
+We use rank-based correlation metrics (i.e., RankIC, RankICIR) instead of error-based ones (e.g., MSE, MAE), as they better capture a factor’s ability to preserve the correct cross-sectional ordering of returns.
+As shown in Table  4, STORM achieves substantial average improvements of 14.690% on SP500 and 15.360% on DJ30, demonstrating its ability to capture more consistent and generalizable signals across diverse market conditions.
+Figure 3. (a) Distribution of counts for codebook embedding indices. (b) Hyperparameter experiment results for different codebook sizes.
+Factor Diversity.
 The diversity of factors is essential in financial investment, as it allows the model to capture the multi-dimensional characteristics of the market, avoiding redundancy or over-reliance on a single factor.
@@
 The results indicates a relatively even distribution of factor usage, highlighting the model’s effectiveness to maintain diversity during the factor selection process.
-Furthermore, we experiment with different codebook sizes on the DJ30 dataset to find the optimal balance between factor diversity and representation quality.
-A larger codebook may lead to sparse factor representations, reducing the effectiveness of factor utilization, while a smaller codebook could result in insufficient factor differentiation and information redundancy.
+Codebook Design.
+Furthermore, we conduct a hyperparameter analysis on the DJ30 dataset to examine the effect of different codebook sizes, aiming to balance factor diversity and representational quality.
+An excessively large codebook may lead to sparse factor representations, limiting the effectiveness of factor utilization, while an overly small codebook could result in insufficient factor differentiation and information redundancy.
 Following standard practices, we evaluate five different codebook sizes, including 128, 256, 512, 1024, and 2048.
-As shown in Figure 3(b), codebook size of 512 achieves optimal factor prediction performance, striking a balance between expressiveness and efficiency.
-Figure 3. (a) Distribution of counts for codebook embedding indices. (b) the RankIC and RankICIR results across different codebook sizes.
-5.3. Downstream Task Performance (for RQ2)
-We use three types of metrics: profit, risk, and risk-adjusted profit to evaluate STORM’s investment performance on downstream tasks, i.e., portfolio management and algorithmic trading.
-For profitability, STORM outperforms all the baselines, with average improvements of 106.10% on APY over the best-performing baseline in the PM task and 43.64% over the best baseline methods on five stocks in the AT task.
-These results underscore the model’s capability to extract meaningful factors, and optimize decision-making for enhanced financial performance.
-For risk resistance, although STORM does not outperform the best baseline methods in terms of MDD and AVO on the INTC and MSFT datasets for the trading task, this can be attributed to the extreme price fluctuations during the COVID-19 event.
-Both stocks experienced a sharp decline of over 30%, followed by a subsequent rebound.
-Given STORM’s strong profitability capability, its trading strategy tends to aggressively capture trends, which naturally leads to larger drawdowns and increased volatility as a trade-off for higher potential returns.
-Moreover, considering the return and risk simultaneously, STORM performs the best on ASR and CR with an average improvement of 58.64% and 166.83% respectively, over the best baseline result in the PM task, and surpasses the best baseline results across three stocks in the AT task.
-Such an outstanding ability to balance return and risk is particularly valuable in real-world investment, where achieving high returns while effectively managing risk is crucial for long-term financial success.
+As shown in Figure 3(b), a size of 512 achieves the best overall factor prediction performance, offering a favorable trade-off between expressiveness and efficiency.
 5.4. Ablation Study (for RQ3)
@@
 Specifically, the absence of either time-series or cross-sectional factors significantly reduces the model’s effectiveness across various tasks.
-In the stock return prediction task, STORM outperforms both simplified models, with an average improvement of 21.73% on the RankIC metric and 24.54% on the RankICIR metric, highlighting the enhanced accuracy achieved by incorporating both time-series and cross-sectional factors.
+In the stock return prediction task, STORM outperforms both simplified models, with an average improvement of 16.50% on the RankIC metric and 23.79% on the RankICIR metric, highlighting the enhanced accuracy achieved by incorporating both time-series and cross-sectional factors.
 In the PM and AT tasks, STORM and two simplified models all performed well.
 Notably, STORM-w/o-TS exhibited the lowest AVO among the three models, although its profitability did not surpass that of STORM.
-These findings suggest that methods like FactorVAE (Duan et al., 2022), which focus solely on cross-sectional factors, may struggle to capture the full complexity of the financial market.
+These findings suggest that methods like FactorVAE (duan2022factorvae), which focus solely on cross-sectional factors, may struggle to capture the full complexity of the financial market.
 Therefore, integrating both time-series and cross-sectional factors is essential for robust and accurate financial modeling.
@@
 Efficiency.
-STORM achieves remarkable efficiency in real-world trading scenarios.
-Unlike high-frequency trading systems requiring ultra-low latency and specialized hardware, our daily trading strategy operates efficiently on standard CPU hardware.
-This is enabled by a compact model architecture characterized by: i) a minimalist design with only a few transformer blocks (see Appendix B), and ii) a lightweight parameterization comprising 15.2 million FP32 parameters on DJ30.
-This design ensures seamless deployment on edge devices while supporting concurrent market analysis and trading.
+STORM runs efficiently in real-world daily trading without the ultra-low latency or specialized hardware needed for high-frequency systems, even on standard CPUs.
+This is enabled by a compact architecture with only a few transformer layers and 15.2M FP32 parameters on DJ30, enabling deployment on edge devices while supporting concurrent market analysis and trading.
 Effectiveness.
-STORM exhibits exceptional performance compared to baseline methods, showcasing versatility across tasks including PM and AT.
-Unlike baselines that often excel in specific tasks but lack generalizability, STORM consistently delivers robust overall results.
-Moreover, STORM can balance returns and risks by emphasizing comprehensive investment performance over individual financial metrics.
-Collectively, these attributes substantiate that the additional computational cost incurred by STORM is well justified by its robustness and adaptability across diverse tasks and datasets.
+STORM delivers consistently strong performance across diverse tasks, including PM and AT, outperforming baselines that often specialize in one domain but fail to generalize.
+It balances return and risk, targeting comprehensive investment performance rather than over-optimizing for a single metric.
+Robustness and versatility across tasks justify the design of our proposed model.
 Generalization.
-STORM demonstrates robust generalizability by achieving strong performance across markets with varying scales and distinct characteristics.
-It excels on both the DJ30, which reflects large-cap industrial blue chips, and the SP500, representing a broader market across diverse industries.
-Furthermore, its flexible encoder-decoder architecture allows the integration of additional layers such as Mamba  (Gu and Dao, 2023; Dao and Gu, 2024) and ensures applicability to diverse trading environments and asset classes.
+STORM generalizes across markets of varying scales and characteristics, excelling on both DJ30 (large-cap industrial blue chips) and SP500 (broad, diversified industries).
+These representative U.S.
+datasets cover concentrated and diversified indices, and their depth makes them a strong proxy for other markets like China or the U.K..
+The flexible encoder-decoder architecture also allows using additional layers like Mamba  (gu2023mamba; mamba2) and ensures applicability to diverse trading environments and asset classes.
 6. Conclusions and Future Work
-In this paper, we propose STORM, a spatio-temporal latent factor model that leverages the dual VQ-VAE framework to learn both time-series and cross-sectional factors.
-STORM ensures the orthogonality and diversity of factors, which are crucial for effective factor selection in financial trading.
-Extensive experiments demonstrate the superiority of STORM over the state-of-the-art methods across a prediction task and two downstream tasks, showcasing its robust performance and effectiveness.
-In the future, we plan to extend STORM by incorporating more side information, e.g., news, to explore the effects of exogenous factors.
-Appendix A Notations
-We provide the main notations in Table A.5 used in the paper.
-Notation Description
-$N$ total number of stocks
-$W$ window size of the observation interval
-$D$ total number of features of data
-$T$ total number of trading days
-$K$ codebook size
-$p_{i,t}$ price of stock $i$ at day $t$
-$\mathbf{x}$ observed data, which contains stock historical price data $\mathbf{p}$ and indicators $\mathbf{d}$
-$\mathbf{x}^{\prime}_{ts}$, $\mathbf{x}^{\prime}_{cs}$, reconstructed observed data of the TS and CS modules
-$\hat{\mathbf{y}}_{i,t}$, $\mathbf{y}_{i,t}$ predicted and true return of stock $i$ at day $t$
-$\mathbf{w}_{t}$ the portfolio weight vector for $N$ stocks at trading day $t$
-$a_{i,t}$ action of day $t$ on the stock $i$
-$s_{t}$ state of day $t$
-$\mathbf{z}_{e}^{ts}(\mathbf{x})$, $\mathbf{z}_{e}^{cs}(\mathbf{x})$ latent embeddings of the TS and CS modules
-$\mathbf{z}_{q}^{ts}(\mathbf{x})$, $\mathbf{z}_{q}^{cs}(\mathbf{x})$ quantized embeddings of the TS and CS modules
-$\mathbf{z}_{e}(\mathbf{x})$ factor embeddings
-$\mathbf{z}$ posterior or latent factors
-$k$ total number of selected stocks in PM strategy
-$d$ the max number of stocks to be sold out
-Table A.5. Notations in the paper.
-Appendix B Implement Details
-We implement all experiments by using PyTorch and conduct the experiments on two 80GB memory NVIDIA RTX H100 GPUs.
-The code is provided in the repository  [^4] https://anonymous.4open.science/r/Storm-2823.
-The input data is organized into batches with dimensions $(B,W,N,D)$, where $B=16$ denotes the batch size.
-$N$ represents the number of stocks, 28 for the DJ30 index and 408 for the SP500 index.
-The dimension $W=64$ corresponds to the number of historical data days, and $D=152$ represents the total number of features, including the OHLC data, technical indicators, and temporal information.
-For the time-series module, we use an 4-day period for each stock as a patch, meaning the patch size is $(4,1,D)$.
-In contrast, for the cross-sectional module, we use $N$ stocks for each day as a patch, so the patch size is $(1,N,D)$.
-For both the encoders and decoders, we employ stacked transformer blocks.
-Each encoder consists of 4 layers of blocks with 4 attention heads, while each decoder consists of 2 layers of blocks with 8 attention heads.
-The size of both codebooks is selected from the set {128, 256, 512, 1024, 2048}, with 512 yielding the best results.
-A consistent embedding dimension of 256 is employed across the encoders, decoders, and codebooks.
-For the optimizer, we use AdamW with a learning rate of $1\times 10^{-4}$ and a weight decay of 0.05.
-The learning rate starts at 0, ramps up to $1\times 10^{-4}$ over 100 epochs with a linear warmup scheduler, and then decays to 0 over 1000 epochs.
-For the loss function balancing, the coefficients for the clip loss, as well as the price reconstruction losses for both the TS and CS modules, are all set to $1\times 10^{-3}$.
-The commitment loss coefficient for the codebooks is 1.0, while the coefficients for the orthogonality loss and diversity loss are both set to 0.1.
-Additionally, the coefficient for the price prediction loss of the factor module, as well as the KL divergence loss between the posterior and prior distributions, is also set to 0.1.
-For downstream tasks, we employ the TopK-Drop rule strategy for portfolio management, using the recommended parameters from Qlib, where $k=5$ and $d=3$.
-For the trading task, we adopt the Proximal Policy Optimization (PPO) RL strategy.
-The CS and TS factors from the pretraining phase are used as the state.
-Notably, we retain all CS factors, which represent market characteristics, while only preserving the TS factors relevant to the currently traded stocks, reflecting stock-specific characteristics.
-The action space has a dimension of 3, corresponding to the choices of buy, hold, and sell.
-Our reward calculation considers value to account for scenarios where remaining cash may not cover the purchase of a new stock.
-Relying solely on price changes without considering position and cash factors is unreasonable.
-Value is defined as $v=p\times m+c$ where $p$ is the current price, $m$ is the position, and $c$ is the remaining cash.
-The reward is calculated as $r=\frac{v_{\text{post}}-v_{\text{pre}}}{v_{\text{pre}}}.$ Unlike compound interest, our method measures the rate of change from the initial value to the current value.
-For both downstream tasks, we start with an initial cash of $1\times 10^{6}$, while also accounting for a transaction cost ratio of $1\times 10^{-4}$.
-For the pretraining prediction tasks and the downstream portfolio management tasks, each set of experiments is conducted 10 times with different random seeds, and the mean and variance are reported.
-For the trading tasks, each set of experiments is conducted 5 times with different random seeds, and the mean of the evaluation metrics is reported.
-Notably, the default parameters provided by the official implementations are used for baseline models, ensuring a consistent and fair comparison across different approaches.
+In this paper, we propose STORM, a spatio-temporal latent factor model leveraging a dual VQ-VAE framework to learn both time-series and cross-sectional factors.
+STORM ensures factor orthogonality and diversity, crucial for effective factor selection in financial trading.
+Unlike traditional factors, STORM learns high-dimensional latent vectors that, though less interpretable, capture richer market structures.
+Extensive experiments demonstrate STORM’s superiority over the state-of-the-art methods across a prediction task and two downstream tasks, showcasing robust performance and effectiveness.
+In the future, we plan to extend STORM by incorporating side information like news, to explore the effects of exogenous factors.
+Acknowledgments
+This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).
+It is also supported, in part, by Jinan-NTU Green Technology Research Institute (GreenTRI).